### v1

# Deep Modularity Networks with Diversity–Preserving Regularization

Abstract
Graph clustering plays a crucial role in graph representation learning but often faces challenges in achieving feature-space diversity. While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they do not explicitly encourage diversity in the feature space among clusters. We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for intra-cluster diversity, and entropy-based for balanced assignments. Our method enhances clustering performance on benchmark datasets, namely Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics, achieving significant improvements in Normalized Mutual Information (NMI), and F1 scores. These results demonstrate the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters, especially in feature-rich datasets.
1 Introduction
Graph clustering is a crucial problem within the graph representation learning domain, and is essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021). Consequently, there has been a surge in methods aimed at enhancing graph clustering performance. Recent advancements have included techniques like Graph Neural Networks (GNNs) (Scarselli et al., 2008), which leverage node features and graph structure for representation learning, often through unsupervised training (Tsitsulin et al., 2023). In this context, graph pooling methods have also become prominent, as they provide a way to coarsen graphs by aggregating nodes into clusters (Cangea et al., 2018). Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly, or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
To address these limitations, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) were introduced, combining spectral modularity maximization (Newman, 2006) with collapse regularization (Tsitsulin et al., 2023) to avoid trivial clustering solutions. While DMoN improves structural separation, it does not explicitly enforce diversity among clusters in the feature space. This limitation can result in structurally distinct but feature-wise homogeneous clusters, reducing effectiveness in applications requiring diverse and meaningfully differentiated clusters.
Present Work
To overcome the limitations of DMoN and enhance feature-space diversity in graph clustering, we propose an extension called Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR). Our method augments the original DMoN optimization objective by introducing additional regularization terms specifically designed to promote diversity among clusters. These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term, which maximizes the variance within each cluster’s assignment probabilities to capture a broader range of node features; and an entropy-based regularization term, which ensures a more balanced and uniform node distribution by maximizing the entropy of cluster assignments.
Our proposed DMoN-DPR method addresses the shortcomings of existing approaches by explicitly enforcing feature-space diversity and ensuring that clusters remain well-separated, all while preventing them from becoming overly homogeneous in their node features. As a result, the clusters produced are both structurally cohesive and feature-wise distinct. Following the evaluation framework in (Tsitsulin et al., 2023), we demonstrate through extensive experiments that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
2 Related Work
Early methods for clustering on graphs often treated node features independently from graph structure. A straightforward example is k-means(features) (Lloyd, 1982), which clusters nodes solely based on attributes and disregards their connectivity. While simple, this approach often overlooks crucial structural relationships. To incorporate at least some graph information, k-means(DeepWalk) (Perozzi et al., 2014) combines node attributes with random-walk-based embeddings. This introduces a basic integration of graph structure but still relies on a two-step process of embedding and clustering. A more advanced alternative is k-means(DGI) (Tsitsulin et al., 2023), which employs node representations learned via Deep Graph Infomax (Veličković et al., 2018) and thereby captures both local and global context. However, because these k-means strategies only merge attributes and structure post hoc—rather than jointly optimizing them—they may lag behind end-to-end trainable methods that unify both graph topology and node features in a single framework.
Moving toward more integrated approaches, Defferrard et al. (2016) introduced one of the first spectral graph convolution methods based on Chebyshev polynomials. Their technique approximates the spectral filters necessary for convolutional operations on graphs, reducing the computational cost of graph Fourier transforms and providing a scalable foundation for subsequent GNN designs. This seminal work illustrated how learned transformations on graph-structured data could drastically improve learning efficiency and accuracy, motivating the community to delve deeper into neural approaches.
Beyond simply learning node-level representations, pooling mechanisms aim to capture hierarchical structures within graphs. Neural Overlapping Community Detection (NOCD) (Shchur and Günnemann, 2019) applies a probabilistic model that directly optimizes the likelihood of reconstructing the graph. Although effective in certain scenarios, NOCD may face scalability challenges and underutilizes node features in complex datasets. One of the earliest neural pooling frameworks, DiffPool (Ying et al., 2018), introduced a soft-clustering strategy that learns cluster assignments in an end-to-end manner. Despite its strong performance, its quadratic time complexity renders it less practical for large-scale graphs (Tsitsulin et al., 2023).
To address computational bottlenecks and overcome potential overfitting, MinCut Pooling (Bianchi et al., 2020) reimagines the clustering problem through a spectral lens, incorporating normalized minCUT objectives with orthogonality constraints. While these constraints can improve cluster quality, they frequently lead to convergence issues. A related simplified variant, Ortho (Bianchi et al., 2020), focuses exclusively on orthogonality regularization, thus discarding many graph-specific objectives. This omission can produce suboptimal results, as the method forgoes some of the structural cues integral to good cluster formation. In another line of work, SAGPool (Lee et al., 2019) introduces a self-attention mechanism that ranks node importance, thereby selecting the most informative nodes during pooling. Although SAGPool effectively combines feature and topology considerations, computing attention coefficients can increase training costs, particularly for large graphs.
Orthogonal to pooling-based approaches, Deep Graph Infomax (DGI) (Velickovic et al., 2019) pioneered an approach that maximizes mutual information between local substructures and a global summary of the graph. This method facilitates robust, label-agnostic node embeddings. The framework was later extended by InfoGraph (Sun et al., 2019), which applies a similar principle at the graph-level, yielding state-of-the-art results for unsupervised and semi-supervised tasks by capturing interactions between subgraphs and a holistic representation.
Most recently, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) have demonstrated a powerful alternative by directly incorporating modularity optimization with collapse regularization in an end-to-end trainable pipeline. DMoN bridges traditional graph community detection and neural feature extraction, achieving strong empirical performance on both synthetic and real-world graphs. By jointly optimizing for high modularity and balancing cluster assignment, DMoN outperforms many existing approaches in terms of normalized mutual information (NMI) and modularity scores, thus setting a new standard for clustering-oriented GNNs.
3 Deep Modularity Networks with Diversity–Preserving Regularization
DMoN tackles the issues of previous techniques by leveraging an optimization objective that combines insights from spectral modularity maximization (Newman, 2006) with a unique regularization term called collapse regularization (Tsitsulin et al., 2023). Specifically, DMoN encodes the cluster assignments, represented as a soft assignment matrix $C$ by using a softmax function over the output of a GNN, allowing differentiation during optimization. For each node, a soft cluster assignment $C$ is computed as follows:
$C=\text{softmax}(\text{GCN}(\tilde{A},X))$ (1)
where GCN is a multi-layer graph convolutional network, $\tilde{A}$ is the normalized adjacency matrix $\tilde{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, and $X$ represents the node features. The objective function of DMoN, $L_{\text{DMoN}}$, combines a modularity term with a collapse regularization term to optimize the clustering. The modularity term measures the quality of the cluster assignments by maximizing the density of intra-cluster edges relative to a null model, while the collapse regularization prevents trivial solutions where all nodes are assigned to the same cluster. The objective function is formulated as:
$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
where $\frac{1}{2m}\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, $\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters. This regularization term encourages balanced cluster assignments, thereby improving the quality of clustering and avoiding degenerate solutions.
Building upon the DMoN framework, we augment the DMoN objective function $L_{\text{DMoN}}$ with three new regularization terms, resulting in the following enhanced objective function:
$L_{\text{DMoN--DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\cdot L_{\text{DPR}}^{\text{Distance}}+W_{\text{var}}\cdot L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\cdot L_{\text{DPR}}^{\text{entropy}}$ (3)
Here, $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$ are hyperparameters that control the influence of each regularization term. The additional terms are designed to promote inter-cluster separation, intra-cluster diversity, and balanced cluster assignments.
Distance-Based Regularization
Inspired by SimCLR (Chen et al., 2020), which showed that contrastive loss promotes well-separated clusters, the distance-based regularization term encourages distinct cluster centroids in feature space. It is defined as:
$L_{\text{DPR}}^{\text{Distance}}=\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\text{ReLU}\left(\epsilon-\|\mu_{i}-\mu_{j}\|_{2}^{2}\right)$ (4)
where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss. By penalizing pairs of clusters whose centroids are closer than $\epsilon$, clusters can be effectively pushed apart in the feature space. By encouraging greater separation between clusters, it enhances inter-cluster diversity and reduces overlap, leading to more distinguishable and meaningful clusters.
Variance-Based Regularization
The variance-based regularization term aims to increase the diversity within each cluster by maximizing the variance of node assignments in each cluster:
$L_{\text{DPR}}^{\text{variance}}=-\frac{1}{k}\sum_{i=1}^{k}\text{Var}(C_{:i})$ (5)
[…]
Entropy-Based Regularization
The entropy-based regularization term promotes a uniform distribution of nodes across clusters, preventing any single cluster from dominating. It is formulated as:
$L_{\text{DPR}}^{\text{entropy}}=-\frac{1}{n}\sum_{v=1}^{n}\sum_{i=1}^{k}C_{vi}\log(C_{vi}+\delta)$ (6)
where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability. By maximizing the entropy of the cluster assignments, this term encourages uncertainty in the assignments, promoting a more balanced distribution of nodes among clusters. This prevents scenarios where one cluster might contain a majority of the nodes, thus enhancing overall diversity and cluster balance.
Therefore, by integrating these diversity-preserving regularization terms into the DMoN objective, we can simultaneously consider graph structure, cluster balance, inter-cluster separation, and intra-cluster diversity.
4 Experiments
Datasets
To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on the widely used Cora, CiteSeer, and PubMed datasets Yang et al. (2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018). The Cora, CiteSeer, and PubMed datasets are citation networks where nodes represent papers and edges denote citation relationships, with node labels corresponding to the topic of each paper. The Coauthor CS and Coauthor Physics datasets are co-authorship graphs, where nodes represent authors and edges indicate collaborations, with labels reflecting the field of research. The key characteristics of these datasets are summarized in Table 1.
Table 1: Dataset characteristics, where ($|V|$) is the number of vertices , ($|E|$) is the number of edges, $|X|$ is the number of features, and $|Y|$ is the number of cluster labels.
Dataset $|V|$ $|E|$ $|X|$ $|Y|$
Cora 2708 5278 1433 7
CiteSeer 3327 4614 3703 6
PubMed 19717 44325 500 3
Coauthor CS 18333 81894 6805 15
Coauthor Phys 34493 247962 8415 5
Evaluation Protocol
Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure. These metrics collectively capture the structural quality of clusters and their alignment with ground-truth labels.
Baseline
We select DMoN as our sole baseline because it consistently achieves state-of-the-art results across multiple datasets and metrics, as shown in prior work (Tsitsulin et al., 2023). It outperforms a wide range of established methods—such as k-means(features) (Lloyd, 1982), k-means(DeepWalk) (Perozzi et al., 2014), k-means(DGI) (Veličković et al., 2018), NOCD (Shchur and Günnemann, 2019), DiffPool (Ying et al., 2018), MinCut Pooling (Bianchi et al., 2020), and Ortho (Bianchi et al., 2020)—each representing distinct approaches that prioritize node attributes, graph structure, or both. Unlike traditional clustering methods like k-means (which lacks graph structure awareness) or embedding-based techniques (which emphasize representation learning over explicit clustering), DMoN offers an end-to-end differentiable framework that directly optimizes modularity and uses collapse regularization to avoid trivial solutions. Furthermore, it exhibits superior scalability and computational efficiency compared to pooling-based approaches such as DiffPool and MinCutPool, making it particularly suited for the large, feature-rich datasets in our experiments. By focusing on feature-space diversity within DMoN’s modularity-driven clustering paradigm, we ensure that our enhancements remain aligned with state-of-the-art structural considerations, providing a clear and meaningful evaluation of the improvements introduced by our proposed modifications.
Implementation Details
Similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by the DMoN/DMoN–DPR pooling layer, as illustrated in Figure 1. Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001. The experiments were run on an Apple M2 Max CPU. The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– will be made publicly available at www.github.com/TBA.
Figure 1: Overview of the DMoN/DMoN–DPR pipeline.
Hyperparameter Tuning
The weighting coefficients $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$, as well as $\epsilon$ are hyperparameters that need to be tuned based on the specific dataset and desired clustering behavior. They control the trade-off between structural modularity and diversity preservation. To optimize these hyperparameters, we conducted evaluations on the Cora, CiteSeer, PubMed, Coauthor CS and Coauthor Physics datasets using 10 random seeds, similar to how it was done in (Tsitsulin et al., 2023). The seeds that resulted in the best performance when using the DMoN pooling layer were selected. For each dataset, we tuned $\epsilon$ (for the distance weight $W_{\text{dist}}$), $W_{\text{var}}$ (variance weight), and $W_{\text{entropy}}$ (entropy weight) independently. To find the best $\epsilon$, the variance and entropy weights were set to zero, and $\epsilon$ was varied from 10 to $10^{-5}$. Similarly, $W_{\text{var}}$ and $W_{\text{entropy}}$ were tuned by fixing the other weights to zero and varying them across {1, 0.1, 0.01, 0.001}. The best weights identified were used to construct various DMoN-DPR models, including DMoN (baseline), DMoN–DPR(D) (distance), DMoN–DPR(V) (variance), DMoN–DPR(E) (entropy), and combinations: DMoN–DPR(DV) (distance and variance), DMoN–DPR(DE) (distance and entropy), DMoN–DPR(VE) (variance and entropy), and DMoN–DPR(DVE) (distance, variance, and entropy). This systematic approach allowed us to assess the impact of each regularization term and their combinations on model performance. Ablation studies are presented in Section A.
5 Results
Table 2: Comparison of clustering methods on three datasets (Cora, Citeseer, Pubmed). Values are in percentage.
Method Cora Citeseer Pubmed
Graph Labels Graph Labels Graph Labels
$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
DMoN 11.00 72.39 43.65 46.24 4.92 75.72 30.26 42.17 8.66 57.03 22.68 45.18
DMoN–DPR(D) 10.23 72.97 45.31 50.00 4.75 75.64 30.61 42.76 8.66 57.02 22.68 45.19
DMoN–DPR(V) 11.64 70.80 44.27 48.18 4.93 75.59 30.46 42.26 8.66 57.02 22.68 45.18
DMoN–DPR(E) 11.00 72.37 43.63 46.25 5.02 75.43 30.52 42.54 8.67 57.02 22.68 45.19
DMoN–DPR(DV) 10.37 72.68 45.41 50.13 4.76 75.51 30.59 42.85 8.66 57.02 22.68 45.19
DMoN–DPR(DE) 10.33 72.88 45.18 49.88 4.86 75.35 30.77 43.27 8.66 57.02 22.67 45.19
DMoN–DPR(VE) 11.64 70.77 44.15 48.09 5.09 75.26 30.86 42.86 8.66 57.01 22.68 45.18
DMoN–DPR(DVE) 10.41 72.65 45.43 50.06 4.85 75.14 30.77 43.42 8.66 57.02 22.67 45.18
Table 3: Comparison of clustering methods on two datasets (Coauthor CS, Coauthor Physics). Values are in percentage.
Method Coauthor CS Coauthor Phys
Graph Labels Graph Labels
$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
DMoN 18.80 72.43 69.45 58.83 13.81 63.35 52.40 46.80
DMoN–DPR(D) 19.01 72.17 70.56 61.09 16.45 56.93 53.99 50.96
DMoN–DPR(V) 18.89 72.41 69.76 59.10 13.45 59.38 54.03 51.89
DMoN–DPR(E) 19.73 70.97 71.34 60.40 13.80 60.04 53.17 50.05
DMoN–DPR(DV) 18.95 72.26 71.33 61.81 13.46 56.85 54.09 56.36
DMoN–DPR(DE) 19.88 70.56 70.84 61.76 14.22 56.10 54.71 56.94
DMoN–DPR(VE) 19.70 71.01 71.45 60.56 12.85 55.72 50.14 54.52
DMoN–DPR(DVE) 19.74 70.68 71.08 62.41 13.46 55.27 54.31 58.86
5.1 Effectiveness of DMoN–DPR in Clustering
To assess the effectiveness of DMoN–DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 seeds achieved with DMoN–DPR to those obtained with DMoN, as done in (Tsitsulin et al., 2023) (full results can be found in Section B). Tables 2 (Cora, CiteSeer, PubMed) and 3 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
Cora, CiteSeer, and PubMed.
On Cora, where the diversity of node features and class labels is relatively high, incorporating the distance-based term (D) consistently enhances performance over the baseline DMoN, especially on label-based metrics (NMI and F1). The combinations that include the distance term—such as DMoN–DPR(DV), DMoN–DPR(DE), and DMoN–DPR(DVE)—not only improve cluster separability but also boost label-based performance, with DMoN–DPR(DV) and DMoN–DPR(DVE) achieving some of the best F1 and NMI scores. The entropy (E) and variance (V) terms alone are less impactful, but when paired with distance, they synergize to produce more balanced and diverse clusters. On CiteSeer, which also has a reasonably complex feature space, the introduction of these regularizations yields moderate gains, with combinations that include the entropy term slightly improving label-based metrics. In contrast, PubMed, which has a less diverse feature space, sees only marginal benefits. Here, the baseline DMoN already performs well, and the added terms provide limited additional advantage.
Coauthor CS and Coauthor Physics.
We observe a similar pattern on the two Coauthor datasets. Both of these datasets have richer feature spaces than Cora, CiteSeer, and PubMed; however, Coauthor Physics stands out with the largest number of features among all the datasets under consideration (8,415). As shown in Table 3, this higher dimensionality helps amplify the benefits of our diversity-preserving regularizations, especially those involving the distance-based term (D). Notably, on Coauthor Physics, the pairwise F1 score jumps from 46.80 (DMoN) to 58.86 (DMoN–DPR(DVE)), indicating that combining distance, variance, and entropy can substantially enhance the separability of clusters and the alignment with true labels.
On Coauthor CS, our proposed regularizations also offer improvements over baseline DMoN, most evident when distance (D) is paired with either variance (V) or entropy (E). For example, F1 increases from 58.83 (DMoN) to 62.41 (DMoN–DPR(DVE)). Although these gains are meaningful, they are somewhat more moderate than those on Coauthor Physics—likely due to differing distributions of class labels and feature dimensions. Overall, these findings suggest that introducing distance-, variance-, and entropy-based regularizations becomes increasingly beneficial as the dimensionality and diversity of node features grow, leading to more pronounced performance boosts on datasets such as Coauthor Physics.
Analyzing the Influence of Regularization on Clustering Performance.
Our experiments show that the impact of the distance-based, variance-based, and entropy-based regularizations depends heavily on both the feature richness and label diversity of each dataset. On Coauthor CS, and Coauthor Physics, which have relatively higher-dimensional or more varied feature spaces, we observe notable gains in NMI and F1 upon adding these diversity-promoting terms—especially the distance-based (D) regularization, which strongly boosts inter-cluster separation. For instance, DMoN–DPR(D) and its variants (e.g., DMoN–DPR(DV), DMoN–DPR(DE), DMoN–DPR(DVE)) yield significant improvements in label alignment on Coauthor Physics, where the F1 score rises by over 10 percentage points compared to baseline DMoN.
Meanwhile, the variance-based (V) term is helpful in ensuring clusters capture a wider range of node features, an effect that is especially apparent on datasets like Coauthor Physics. Although the single V-only variant often yields moderate gains on datasets such as Cora, it generally performs best when combined with the distance or entropy regularizers. The entropy-based (E) term aims to produce more balanced cluster distributions, proving modestly beneficial on datasets like CiteSeer and effective on Coauthor CS and Coauthor Physics; however, on PubMed, which has lower feature dimensionality, we see little to no improvement (the numbers remain nearly unchanged from DMoN). This indicates that entropy-based balancing may not substantially help in more homogeneous datasets.
Moreover, combining all three terms (DMoN–DPR(DVE)) often yields the highest F1 scores across multiple datasets (e.g., a leap from 58.83 to 62.41 on Coauthor CS and 46.80 to 58.86 on Coauthor Physics). These results suggest that multiple regularizers can work in synergy, but the exact benefits depend on a dataset’s inherent complexity. Overall, the findings highlight the importance of tailoring regularization strategies to both the graph structure and feature diversity of the data to achieve the most effective clustering outcomes.
Trade-Offs Between Structural and Label-Based Metrics
Our results highlight a trade-off between metrics that prioritize graph structure (e.g., Modularity ($Q$) and Conductance ($C$)) and those that evaluate label alignment (e.g., NMI and F1). This trade-off emerges because diversity-preserving regularization terms influence both the feature and label spaces of clusters, leading to distinct outcomes depending on the dataset characteristics and application needs.
For datasets with rich feature spaces (e.g., Coauthor CS and Coauthor Physics), the diversity-enhancing terms significantly boost label-based metrics like NMI and F1. These gains stem from the formation of more homogeneous clusters that align closely with ground-truth labels. However, this improvement in label alignment can come at the expense of structural metrics such as modularity, which reward densely connected clusters in the graph topology. For instance, while DMoN–DPR(DVE) achieves the highest F1 scores on Coauthor Physics, it exhibits a decrease in modularity compared to baseline DMoN.
Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest. In such cases, the original DMoN, which focuses on structural cohesion, remains a competitive choice for purely unsupervised clustering tasks.
Ultimately, the decision to prioritize label alignment or structural metrics should be informed by the application context. For tasks involving semi-supervised learning or node classification, where accurate label recovery is essential, DMoN–DPR offers clear advantages. In contrast, for community detection or unsupervised settings where graph structure is paramount, the baseline DMoN may be preferable. This balance between competing objectives underscores the need to tailor clustering approaches to the specific demands of the dataset and task at hand.
Figure 2: Visualization of clusters (circles) vs. ground truth labels (crosses) formed by DMoN on (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics, and by DMoN–DPR on (f) Cora, (g) CiteSeer, (h) PubMed, (i) Coauthor CS, and (j) Coauthor Physics datasets.
5.2 Visualization of Clusters
The clustering results visualized in Figure 1 demonstrate that DMoN–DPR generally outperforms DMoN across the benchmark datasets Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics. For Cora, DMoN–DPR produces tighter, more well-separated clusters with improved alignment to the ground truth labels, contrasting with the more scattered and less defined clusters formed by DMoN. Similarly, for CiteSeer, DMoN–DPR exhibits better cluster compactness and reduced overlap, though the gains are somewhat less pronounced than in Cora. On the other hand, PubMed remains challenging for both methods, resulting in notable overlap and scattered clusters; yet, DMoN–DPR still provides slightly tighter groupings. Across these three datasets, the diversity-preserving regularization in DMoN–DPR effectively enhances both inter-cluster distinctiveness and intra-cluster variability, producing more meaningful and interpretable partitions. Turning to the Coauthor CS and Coauthor Physics datasets, a similar trend emerges. While DMoN adequately captures overarching modular structures, it can occasionally struggle to separate nuanced subgroups. In contrast, DMoN–DPR consistently yields more cohesive and better-separated clusters, offering clearer decision boundaries and higher alignment with the true labels. This improvement is particularly notable in Coauthor CS, where DMoN–DPR visibly reduces overlap and sharpens cluster delineations. Although the inherently complex Coauthor Physics dataset can induce some overlap in both methods, DMoN–DPR still produces more robust clusters overall. These findings underscore the effectiveness of promoting diversity within hidden representations, as DMoN–DPR not only excels in forming compact and accurate clusters for datasets with greater feature variability but also remains competitive on more challenging datasets.
5.3 Execution Times and Runtime Analysis
Table 4 reports the execution times (in seconds) of DMoN and DMoN–DPR on five commonly used graph datasets. All experiments were conducted on a CPU to ensure consistent runtime measurement without GPU scheduling effects.
Table 4: Execution times (in seconds) on CPU for DMoN and DMoN–DPR.
Dataset DMoN DMoN–DPR
Cora 32 35
CiteSeer 76 71
PubMed 310 300
Coauthor CS 722 751
Coauthor Physics 2107 1946
Although DMoN–DPR includes additional regularization terms for Distance-based, Entropy-based, and Variance-based constraints, its runtime is sometimes comparable to or even marginally lower than plain DMoN. The reason is that the most time-consuming part of both methods is dominated by adjacency-based operations (e.g., matrix multiplications with the graph Laplacian or adjacency matrix), which scale on the order of $\mathcal{O}(|E|\,k)$ or $\mathcal{O}(N^{2}k)$ for large, dense graphs (where $N$ is the number of nodes, $E$ is the set of edges, and $k$ is the number of clusters). In contrast, the extra DPR calculations—computing entropy across nodes, the variance of cluster assignments, or centroid distances—represent only $\mathcal{O}(N\,k+k^{2}F)$ overhead (with $F$ being the feature dimension) and are typically negligible next to the larger matrix multiplications.
Furthermore, running on CPU can introduce variability from scheduling, caching, and background processes, so small differences in the recorded times may not be statistically meaningful. Thus, although DMoN–DPR performs slightly more computations, the extra cost is overshadowed by the main bottleneck operations, resulting in comparable or even lower measured runtimes in practice.
5.4 Limitations and Future Directions
Despite the improvements introduced by our diversity–preserving regularization, DMoN–DPR has a few notable limitations. First, tuning the regularization weights for distance, variance, and entropy remains a manual process. Determining the best values for these hyperparameters can be time-consuming and sensitive to each dataset’s characteristics. An interesting direction for future work is to develop an automated or adaptive scheme that learns these weights during training, making the method more robust and less reliant on exhaustive search.
Second, while DMoN–DPR yields substantial gains on datasets with a large or diverse feature space, the improvements are often marginal for datasets with lower feature dimensionality (e.g., PubMed). Nonetheless, our visualization in Figure 2 suggests that even in these cases, DMoN–DPR tends to produce slightly more coherent cluster structures compared to the baseline, reflecting the benefit of promoting diversity even when numerical metrics (e.g., NMI, F1) show limited change. Future research could investigate whether alternative or more specialized regularizers can drive greater numerical improvements on lower-dimensional datasets.
6 Conclusion
In this work, we presented DMoN–DPR, an enhanced version of Deep Modularity Networks (DMoN) that incorporates diversity-preserving regularizations to enrich the clustering objective. By introducing distance-, entropy-, and variance-based penalties alongside the original modularity and collapse regularization terms, our approach promotes both inter- and intra-cluster diversity. Our empirical results on benchmark datasets indicate that DMoN–DPR significantly improves alignment with ground-truth labels, particularly in feature-rich datasets. While structural metrics such as modularity and conductance remained competitive, the additional diversity constraints contributed to more interpretable and semantically meaningful cluster formations. Overall, DMoN–DPR achieves a balance between interpretability and structural integrity, making it a powerful solution for graph clustering tasks that demand both topological cohesion and semantic differentiation, marking a significant advancement in feature-aware graph clustering.
Appendix A Ablation Study
A.1 Varying the Epsilon
Figure 3 depicts the ablation study done to find the best value of $\epsilon$ associated with $W_{\text{dist}}$ by setting $W_{\text{var}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $\epsilon$ values were found to be $0.0001$, $0.0001$, $0.001$, $1.0$, and $10.0$ respectively.
Figure 3: Finding the best value of epsilon ($\epsilon$) by setting $W_{\text{dist}}$ to 1 and $W_{\text{var}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
A.2 Varying the Variance Weight
Figure 4 depicts the ablation study done to find the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be $1$/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
Figure 4: Finding the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
A.3 Varying the Entropy Weight
Figure 5 depicts the ablation study done to find the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ weights to 0. For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{entropy}}$ values were found to be $0.001$, $0.01$, $0.001$, $0.1$ and $0.1$ respectively.
Figure 5: Finding the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
Appendix B Full Results
The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets. The numbers are all in percentage (%).
B.1 Cora
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
Table 5: Results obtained by using DMoN pooling on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 10.25 73.93 48.71 55.38
550 DMoN 9.95 71.77 40.62 42.17
243 DMoN 10.06 73.74 47.62 53.20
16 DMoN 12.13 70.62 44.77 46.61
716 DMoN 11.92 70.40 37.87 39.65
383 DMoN 11.20 71.80 44.40 52.46
277 DMoN 11.03 72.50 37.07 36.07
274 DMoN 10.12 74.14 45.30 44.74
188 DMoN 11.48 73.36 46.79 49.47
796 DMoN 11.86 71.67 43.30 42.69
Mean 11.00 72.39 43.65 46.24
Variance 0.01 0.02 0.16 0.40
Table 6: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, and distance weight of 1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 10.50 73.78 51.01 57.45
550 DMoN–DPR (D) 10.06 72.20 41.49 44.43
243 DMoN–DPR (D) 9.49 73.58 47.80 54.94
16 DMoN–DPR (D) 10.55 72.20 46.22 49.19
716 DMoN–DPR (D) 10.38 71.10 39.29 43.43
383 DMoN–DPR (D) 9.49 74.11 48.02 54.77
277 DMoN–DPR (D) 10.65 72.45 40.12 42.77
274 DMoN–DPR (D) 9.47 74.41 47.64 52.47
188 DMoN–DPR (D) 10.44 73.75 48.41 54.78
796 DMoN–DPR (D) 11.29 72.13 43.05 45.75
Mean 10.23 72.97 45.31 50.00
Variance 0.00 0.01 0.16 0.31
Table 7: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR 10.99 72.41 51.86 58.92
550 DMoN–DPR (V) 11.10 70.31 40.04 43.09
243 DMoN–DPR (V) 11.75 71.02 47.47 54.54
16 DMoN–DPR (V) 10.99 70.18 45.82 48.81
716 DMoN–DPR (V) 11.08 69.84 40.82 41.88
383 DMoN–DPR (V) 12.13 69.37 42.89 51.12
277 DMoN–DPR (V) 11.42 70.91 37.43 37.64
274 DMoN–DPR (V) 11.05 72.61 46.97 48.40
188 DMoN–DPR (V) 13.11 70.74 45.71 52.26
796 DMoN–DPR (V) 12.79 70.56 43.70 45.13
Mean 11.64 70.80 44.27 48.18
Variance 0.01 0.01 0.18 0.41
Table 8: Results obtained by using DMoN–DPR pooling, using the best entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 10.25 73.81 49.10 55.86
550 DMoN–DPR (E) 10.06 71.70 40.49 42.06
243 DMoN–DPR (E) 9.91 73.82 48.60 53.72
16 DMoN–DPR (E) 12.32 70.46 44.95 46.88
716 DMoN–DPR (E) 11.86 70.43 37.72 39.66
383 DMoN–DPR (E) 11.01 71.93 44.49 52.83
277 DMoN–DPR (E) 11.16 72.37 37.38 35.98
274 DMoN–DPR (E) 10.17 74.20 44.13 44.02
188 DMoN–DPR (E) 11.41 73.37 46.28 48.86
796 DMoN–DPR (E) 11.88 71.65 43.17 42.63
Mean 11.00 72.37 43.63 46.25
Variance 0.01 0.02 0.16 0.43
Table 9: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 10.25 73.42 51.71 58.37
550 DMoN–DPR (DV) 10.17 72.08 41.28 44.08
243 DMoN–DPR (DV) 9.66 73.18 48.08 55.94
16 DMoN–DPR (DV) 10.72 71.69 46.80 49.51
716 DMoN–DPR (DV) 10.52 70.97 39.46 43.32
383 DMoN–DPR (DV) 9.87 73.82 48.11 54.87
277 DMoN–DPR (DV) 10.95 71.90 39.01 42.43
274 DMoN–DPR (DV) 9.61 74.13 48.04 53.26
188 DMoN–DPR (DV) 10.57 73.51 48.47 54.55
796 DMoN–DPR (DV) 11.39 72.06 43.12 44.93
Mean 10.37 72.68 45.41 50.13
Variance 0.00 0.01 0.19 0.36
Table 10: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 10.63 73.82 51.30 57.10
550 DMoN–DPR (DE) 10.10 72.15 41.06 44.04
243 DMoN–DPR (DE) 9.68 73.25 47.62 55.30
16 DMoN–DPR (DE) 10.61 72.13 46.23 49.09
716 DMoN–DPR (DE) 10.63 70.90 39.37 43.28
383 DMoN–DPR (DE) 9.53 74.04 47.93 54.43
277 DMoN–DPR (DE) 10.89 72.25 39.14 42.53
274 DMoN–DPR (DE) 9.49 74.39 48.31 52.84
188 DMoN–DPR (DE) 10.52 73.68 48.16 54.61
796 DMoN–DPR (DE) 11.18 72.21 42.71 45.56
Mean 10.33 72.88 45.18 49.88
Variance 0.00 0.01 0.18 0.32
Table 11: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 10.86 72.34 51.43 58.35
550 DMoN–DPR (VE) 11.10 70.30 40.10 43.20
243 DMoN–DPR (VE) 11.71 71.05 47.46 54.42
16 DMoN–DPR (VE) 11.06 70.19 45.89 48.85
716 DMoN–DPR (VE) 11.01 69.83 40.77 41.82
383 DMoN–DPR (VE) 12.14 69.28 42.65 50.92
277 DMoN–DPR (VE) 11.46 70.98 37.40 37.53
274 DMoN–DPR (VE) 11.14 72.47 46.72 48.43
188 DMoN–DPR (VE) 13.11 70.78 45.36 52.19
796 DMoN–DPR (VE) 12.85 70.48 43.74 45.16
Mean 11.64 70.77 44.15 48.09
Variance 0.01 0.01 0.17 0.39
Table 12: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, variance weight of 0.1, and entropy weight of 0.001 on the Cora dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 10.06 73.55 51.62 58.48
550 DMoN–DPR (DVE) 10.16 72.04 41.47 44.30
243 DMoN–DPR (DVE) 9.87 72.96 47.86 55.57
16 DMoN–DPR (DVE) 10.76 71.64 46.63 49.25
716 DMoN–DPR (DVE) 10.72 70.81 39.64 43.32
383 DMoN–DPR (DVE) 9.93 73.75 48.14 54.86
277 DMoN–DPR (DVE) 10.89 71.99 39.10 42.50
274 DMoN–DPR (DVE) 9.83 73.94 47.81 52.85
188 DMoN–DPR (DVE) 10.42 73.73 48.53 54.50
796 DMoN–DPR (DVE) 11.41 72.05 43.51 44.96
Mean 10.41 72.65 45.43 50.06
Variance 0.00 0.01 0.18 0.35
B.2 CiteSeer
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
Table 13: Results obtained by using DMoN pooling on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 6.09 74.10 22.56 26.60
550 DMoN 4.31 75.22 25.92 41.43
243 DMoN 4.90 78.26 34.20 46.62
16 DMoN 5.87 75.48 28.89 38.28
716 DMoN 3.60 76.57 38.40 52.01
383 DMoN 4.94 75.34 26.50 39.64
277 DMoN 6.50 75.67 28.87 37.96
274 DMoN 4.35 74.93 32.36 47.53
188 DMoN 4.50 76.41 34.86 46.98
796 DMoN 4.13 75.24 30.07 44.67
Mean 4.92 75.72 30.26 42.17
Variance 0.01 0.01 0.23 0.51
Table 14: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, and distance weight of 1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 6.30 74.07 22.96 27.06
550 DMoN–DPR (D) 4.11 75.45 26.52 42.31
243 DMoN–DPR (D) 3.98 76.17 36.63 51.64
16 DMoN–DPR (D) 5.36 76.07 29.32 42.93
716 DMoN–DPR (D) 4.17 76.07 39.05 51.65
383 DMoN–DPR (D) 5.18 75.60 26.83 37.54
277 DMoN–DPR (D) 5.16 75.30 29.25 40.17
274 DMoN–DPR (D) 4.28 75.19 32.02 46.99
188 DMoN–DPR (D) 4.70 76.60 34.34 45.97
796 DMoN–DPR (D) 4.24 75.92 29.15 41.29
Mean 4.75 75.64 30.61 42.76
Variance 0.01 0.00 0.24 0.52
Table 15: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 6.33 73.98 23.33 26.50
550 DMoN–DPR (V) 4.17 75.20 26.64 42.18
243 DMoN–DPR (V) 4.99 78.15 34.08 46.34
16 DMoN–DPR (V) 6.02 75.21 29.38 38.80
716 DMoN–DPR (V) 3.71 76.45 38.70 52.38
383 DMoN–DPR (V) 5.16 74.60 26.80 39.65
277 DMoN–DPR (V) 5.84 75.26 29.13 39.42
274 DMoN–DPR (V) 4.42 74.79 31.87 46.92
188 DMoN–DPR (V) 4.33 76.39 34.76 47.14
796 DMoN–DPR (V) 4.33 75.87 29.92 43.30
Mean 4.93 75.59 30.46 42.26
Variance 0.01 0.01 0.20 0.49
Table 16: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 6.59 74.00 22.17 27.07
550 DMoN–DPR (E) 4.42 75.06 26.32 41.61
243 DMoN–DPR (E) 5.14 77.95 34.48 46.79
16 DMoN–DPR (E) 6.15 75.09 29.46 38.54
716 DMoN–DPR (E) 3.67 76.36 39.12 52.80
383 DMoN–DPR (E) 5.32 74.52 27.05 39.70
277 DMoN–DPR (E) 5.89 75.27 29.14 39.34
274 DMoN–DPR (E) 4.35 74.85 32.18 47.28
188 DMoN–DPR (E) 4.55 76.16 34.69 47.23
796 DMoN–DPR (E) 4.11 75.02 30.58 45.06
Mean 5.02 75.43 30.52 42.54
Variance 0.01 0.01 0.24 0.50
Table 17: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 6.33 73.81 23.64 27.24
550 DMoN–DPR (DV) 4.09 75.48 26.65 42.08
243 DMoN–DPR (DV) 3.89 76.17 35.67 51.00
16 DMoN–DPR (DV) 5.71 75.73 29.33 42.91
716 DMoN–DPR (DV) 4.11 76.03 38.91 51.73
383 DMoN–DPR (DV) 5.16 75.31 27.06 38.22
277 DMoN–DPR (DV) 5.01 75.32 29.81 40.68
274 DMoN–DPR (DV) 4.35 74.92 32.29 47.22
188 DMoN–DPR (DV) 4.50 76.36 34.12 46.47
796 DMoN–DPR (DV) 4.44 75.93 28.46 40.94
Mean 4.76 75.51 30.59 42.85
Variance 0.01 0.01 0.21 0.50
Table 18: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 6.48 73.71 23.56 27.55
550 DMoN–DPR (DE) 4.31 75.24 26.29 41.96
243 DMoN–DPR (DE) 3.95 76.11 36.10 51.46
16 DMoN–DPR (DE) 5.76 75.68 29.78 43.37
716 DMoN–DPR (DE) 4.15 76.00 39.03 51.88
383 DMoN–DPR (DE) 5.43 75.09 27.51 38.72
277 DMoN–DPR (DE) 5.10 75.25 29.91 40.77
274 DMoN–DPR (DE) 4.42 74.89 31.85 46.74
188 DMoN–DPR (DE) 4.75 76.28 34.05 46.33
796 DMoN–DPR (DE) 4.26 75.21 29.63 43.94
Mean 4.86 75.35 30.77 43.27
Variance 0.01 0.01 0.22 0.49
Table 19: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and best entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 6.50 73.55 22.91 27.80
550 DMoN–DPR (VE) 4.59 74.88 26.41 41.35
243 DMoN–DPR (VE) 5.25 77.84 34.60 47.05
16 DMoN–DPR (VE) 5.95 76.05 31.65 40.71
716 DMoN–DPR (VE) 3.84 76.18 39.28 52.93
383 DMoN–DPR (VE) 5.40 74.12 26.44 39.40
277 DMoN–DPR (VE) 5.91 75.11 29.14 39.27
274 DMoN–DPR (VE) 4.55 74.52 32.02 47.12
188 DMoN–DPR (VE) 4.70 75.85 34.92 47.37
796 DMoN–DPR (VE) 4.24 74.50 31.26 45.61
Mean 5.09 75.26 30.86 42.86
Variance 0.01 0.02 0.23 0.47
Table 20: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, variance weight of 0.1, and entropy weight of 0.01 on the CiteSeer dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 6.39 73.36 23.60 28.45
550 DMoN–DPR (DVE) 4.35 75.08 26.35 41.77
243 DMoN–DPR (DVE) 3.93 75.80 36.58 52.14
16 DMoN–DPR (DVE) 5.38 76.03 28.50 41.56
716 DMoN–DPR (DVE) 4.22 75.83 38.91 51.92
383 DMoN–DPR (DVE) 5.34 74.92 28.02 39.07
277 DMoN–DPR (DVE) 5.34 75.06 29.87 40.58
274 DMoN–DPR (DVE) 4.46 74.65 31.76 46.98
188 DMoN–DPR (DVE) 4.68 75.90 34.46 47.48
796 DMoN–DPR (DVE) 4.44 74.79 29.69 44.27
Mean 4.85 75.14 30.77 43.42
Variance 0.01 0.01 0.22 0.48
B.3 PubMed
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.
Table 21: Results obtained by using DMoN pooling on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 10.64 55.82 21.92 39.57
550 DMoN 7.91 58.30 26.08 48.88
243 DMoN 7.87 57.66 21.90 44.67
16 DMoN 8.02 57.40 21.90 44.83
716 DMoN 8.02 57.35 21.98 45.16
383 DMoN 12.56 53.65 26.03 49.14
277 DMoN 7.78 57.61 21.73 44.90
274 DMoN 7.94 57.44 21.59 44.78
188 DMoN 7.90 57.50 21.86 45.03
796 DMoN 7.93 57.53 21.76 44.82
Mean 8.66 57.03 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 22: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001 and distance weight of 1 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 10.63 55.83 21.92 39.59
550 DMoN–DPR (D) 7.94 58.27 26.16 48.94
243 DMoN–DPR (D) 7.88 57.65 21.93 44.67
16 DMoN–DPR (D) 8.00 57.41 21.80 44.75
716 DMoN–DPR (D) 8.07 57.30 21.97 45.11
383 DMoN–DPR (D) 12.53 53.69 25.84 49.03
277 DMoN–DPR (D) 7.79 57.60 21.73 44.91
274 DMoN–DPR (D) 7.96 57.43 21.76 44.94
188 DMoN–DPR (D) 7.91 57.49 21.87 45.05
796 DMoN–DPR (D) 7.92 57.54 21.84 44.89
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 23: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 10.64 55.82 21.91 39.57
550 DMoN–DPR (V) 7.92 58.29 26.08 48.88
243 DMoN–DPR (V) 7.88 57.65 21.91 44.68
16 DMoN–DPR (V) 8.01 57.40 21.92 44.86
716 DMoN–DPR (V) 8.02 57.35 21.97 45.16
383 DMoN–DPR (V) 12.56 53.64 26.03 49.14
277 DMoN–DPR (V) 7.78 57.61 21.75 44.89
274 DMoN–DPR (V) 7.94 57.44 21.59 44.78
188 DMoN–DPR (V) 7.91 57.49 21.86 45.04
796 DMoN–DPR (V) 7.92 57.54 21.78 44.83
Mean 8.66 57.02 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 24: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 10.64 55.82 21.91 39.59
550 DMoN–DPR (E) 7.92 58.29 26.08 48.88
243 DMoN–DPR (E) 7.88 57.65 21.90 44.67
16 DMoN–DPR (E) 8.01 57.40 21.90 44.83
716 DMoN–DPR (E) 8.03 57.34 21.96 45.13
383 DMoN–DPR (E) 12.55 53.65 26.04 49.15
277 DMoN–DPR (E) 7.78 57.61 21.75 44.91
274 DMoN–DPR (E) 7.96 57.42 21.63 44.80
188 DMoN–DPR (E) 7.93 57.47 21.88 45.06
796 DMoN–DPR (E) 7.95 57.52 21.79 44.85
Mean 8.67 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 25: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and variance weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 10.63 55.84 21.89 39.58
550 DMoN–DPR (DV) 7.93 58.27 26.15 48.92
243 DMoN–DPR (DV) 7.89 57.64 21.94 44.67
16 DMoN–DPR (DV) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DV) 8.06 57.31 22.00 45.11
383 DMoN–DPR (DV) 12.52 53.69 25.81 49.01
277 DMoN–DPR (DV) 7.79 57.59 21.72 44.91
274 DMoN–DPR (DV) 7.93 57.46 21.73 44.92
188 DMoN–DPR (DV) 7.91 57.48 21.83 45.03
796 DMoN–DPR (DV) 7.93 57.53 21.86 44.93
Mean 8.66 57.02 22.68 45.19
Variance 0.03 0.02 0.03 0.07
Table 26: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 10.60 55.87 21.90 39.66
550 DMoN–DPR (DE) 7.94 58.27 26.15 48.91
243 DMoN–DPR (DE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DE) 7.97 57.43 21.83 44.78
716 DMoN–DPR (DE) 8.08 57.29 22.01 45.14
383 DMoN–DPR (DE) 12.54 53.68 25.80 49.00
277 DMoN–DPR (DE) 7.79 57.59 21.71 44.90
274 DMoN–DPR (DE) 7.97 57.42 21.75 44.94
188 DMoN–DPR (DE) 7.93 57.47 21.86 45.01
796 DMoN–DPR (DE) 7.92 57.54 21.78 44.87
Mean 8.66 57.02 22.67 45.19
Variance 0.03 0.02 0.03 0.07
Table 27: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 10.64 55.83 21.89 39.57
550 DMoN–DPR (VE) 7.92 58.29 26.09 48.88
243 DMoN–DPR (VE) 7.87 57.65 21.89 44.67
16 DMoN–DPR (VE) 8.01 57.40 21.91 44.85
716 DMoN–DPR (VE) 8.03 57.33 21.96 45.13
383 DMoN–DPR (VE) 12.57 53.63 26.02 49.14
277 DMoN–DPR (VE) 7.78 57.60 21.74 44.91
274 DMoN–DPR (VE) 7.96 57.42 21.61 44.81
188 DMoN–DPR (VE) 7.92 57.47 21.90 45.03
796 DMoN–DPR (VE) 7.94 57.52 21.77 44.84
Mean 8.66 57.01 22.68 45.18
Variance 0.03 0.02 0.03 0.07
Table 28: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 10.59 55.87 21.90 39.65
550 DMoN–DPR (DVE) 7.93 58.27 26.14 48.91
243 DMoN–DPR (DVE) 7.89 57.64 21.93 44.67
16 DMoN–DPR (DVE) 7.97 57.42 21.84 44.78
716 DMoN–DPR (DVE) 8.08 57.29 21.91 45.06
383 DMoN–DPR (DVE) 12.51 53.69 25.79 48.98
277 DMoN–DPR (DVE) 7.79 57.58 21.70 44.88
274 DMoN–DPR (DVE) 7.94 57.45 21.74 44.94
188 DMoN–DPR (DVE) 7.94 57.45 21.89 45.04
796 DMoN–DPR (DVE) 7.92 57.54 21.81 44.89
Mean 8.66 57.02 22.67 45.18
Variance 0.03 0.02 0.03 0.07
B.4 Coauthor CS
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
Table 29: Results obtained by using DMoN pooling on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 19.07 72.23 68.48 58.27
550 DMoN 19.14 71.64 66.78 55.19
243 DMoN 18.11 72.88 70.78 60.88
16 DMoN 18.28 73.04 71.13 62.23
716 DMoN 18.41 72.63 69.20 59.44
383 DMoN 18.45 72.43 69.96 59.60
277 DMoN 19.53 72.25 67.51 54.79
274 DMoN 20.20 71.69 69.38 58.09
188 DMoN 17.89 73.05 71.52 60.61
796 DMoN 18.89 72.50 69.77 59.18
Mean 18.80 72.43 69.45 58.83
Variance 0.01 0.00 0.02 0.06
Table 30: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, and distance weight of 1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR(D) 18.88 72.53 72.11 62.71
550 DMoN–DPR(D) 20.80 71.19 68.18 57.76
243 DMoN–DPR(D) 19.36 71.83 71.19 62.87
16 DMoN–DPR(D) 18.99 72.42 71.12 63.30
716 DMoN–DPR(D) 17.85 73.08 70.14 56.54
383 DMoN–DPR(D) 19.24 71.60 69.23 60.49
277 DMoN–DPR(D) 18.20 73.02 72.35 64.04
274 DMoN–DPR(D) 18.86 72.11 70.97 61.56
188 DMoN–DPR(D) 19.85 71.18 68.61 59.59
796 DMoN–DPR(D) 18.02 72.72 71.70 62.08
Mean 19.01 72.17 70.56 61.09
Variance 0.01 0.00 0.02 0.06
Table 31: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 18.87 72.35 68.79 58.77
550 DMoN–DPR (V) 19.03 71.68 67.07 56.15
243 DMoN–DPR (V) 18.27 72.74 71.13 61.15
16 DMoN–DPR (V) 18.28 72.94 71.38 62.48
716 DMoN–DPR (V) 19.77 71.86 68.11 57.05
383 DMoN–DPR (V) 18.53 72.22 70.15 60.21
277 DMoN–DPR (V) 19.78 72.14 67.97 55.30
274 DMoN–DPR (V) 19.57 72.49 70.45 59.36
188 DMoN–DPR (V) 18.24 72.89 71.02 59.87
796 DMoN–DPR (V) 18.52 72.76 71.57 60.67
Mean 18.89 72.41 69.76 59.10
Variance 0.00 0.00 0.03 0.05
Table 32: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 18.95 71.61 72.94 59.97
550 DMoN–DPR (E) 19.87 71.23 70.05 56.24
243 DMoN–DPR (E) 20.13 70.45 71.39 60.47
16 DMoN–DPR (E) 19.44 71.06 72.47 64.02
716 DMoN–DPR (E) 19.11 71.28 73.55 64.42
383 DMoN–DPR (E) 20.09 70.20 67.61 54.84
277 DMoN–DPR (E) 21.27 69.71 68.97 56.84
274 DMoN–DPR (E) 20.01 71.45 73.13 62.46
188 DMoN–DPR (E) 18.70 71.68 72.18 63.95
796 DMoN–DPR (E) 19.74 71.04 71.06 60.77
Mean 19.73 70.97 71.34 60.40
Variance 0.01 0.00 0.04 0.12
Table 33: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and variance weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 18.76 72.48 72.77 63.41
550 DMoN–DPR (DV) 19.71 71.95 70.36 61.52
243 DMoN–DPR (DV) 19.10 72.09 71.75 63.65
16 DMoN–DPR (DV) 18.57 72.77 73.03 63.98
716 DMoN–DPR (DV) 19.02 72.54 70.44 58.08
383 DMoN–DPR (DV) 18.57 72.33 72.29 63.93
277 DMoN–DPR (DV) 17.94 73.14 72.46 63.13
274 DMoN–DPR (DV) 19.41 71.73 70.41 60.45
188 DMoN–DPR (DV) 20.33 71.18 69.03 58.99
796 DMoN–DPR (DV) 18.06 72.38 70.79 60.99
Mean 18.95 72.26 71.33 61.81
Variance 0.01 0.00 0.02 0.05
Table 34: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 19.82 70.82 71.87 64.34
550 DMoN–DPR (DE) 19.72 70.51 70.06 61.80
243 DMoN–DPR (DE) 19.74 70.31 70.33 61.84
16 DMoN–DPR (DE) 19.83 71.73 74.31 63.08
716 DMoN–DPR (DE) 18.97 71.30 71.50 62.63
383 DMoN–DPR (DE) 19.97 70.74 72.47 65.26
277 DMoN–DPR (DE) 19.50 71.15 72.02 60.54
274 DMoN–DPR (DE) 19.42 69.84 68.51 59.93
188 DMoN–DPR (DE) 21.62 69.25 67.61 56.77
796 DMoN–DPR (DE) 20.19 69.93 69.69 61.41
Mean 19.88 70.56 70.84 61.76
Variance 0.00 0.01 0.04 0.06
Table 35: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 19.08 71.49 73.11 60.17
550 DMoN–DPR (VE) 19.95 71.16 70.08 56.20
243 DMoN–DPR (VE) 20.21 70.35 71.26 60.33
16 DMoN–DPR (VE) 19.62 70.89 72.13 63.50
716 DMoN–DPR (VE) 19.16 71.25 73.42 64.30
383 DMoN–DPR (VE) 20.23 70.06 67.62 54.87
277 DMoN–DPR (VE) 20.14 70.81 70.43 59.06
274 DMoN–DPR (VE) 20.04 71.40 73.23 62.31
188 DMoN–DPR (VE) 18.72 71.67 72.17 63.93
796 DMoN–DPR (VE) 19.80 70.97 71.09 60.92
Mean 19.70 71.01 71.45 60.56
Variance 0.00 0.00 0.03 0.10
Table 36: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 19.32 71.13 72.76 65.12
550 DMoN–DPR (DVE) 19.75 70.50 70.00 61.78
243 DMoN–DPR (DVE) 19.97 70.19 70.40 61.86
16 DMoN–DPR (DVE) 18.87 72.23 74.23 65.17
716 DMoN–DPR (DVE) 19.03 71.26 71.56 62.76
383 DMoN–DPR (DVE) 20.06 70.65 72.47 65.18
277 DMoN–DPR (DVE) 19.06 72.07 73.97 64.36
274 DMoN–DPR (DVE) 19.41 69.81 68.57 60.03
188 DMoN–DPR (DVE) 21.66 69.20 67.65 56.86
796 DMoN–DPR (DVE) 20.30 69.78 69.21 60.98
Mean 19.74 70.68 71.08 62.41
Variance 0.01 0.01 0.05 0.07
B.5 Coauthor Physics
The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.
Table 37: Results obtained by using DMoN pooling on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN 12.61 60.14 44.16 46.45
550 DMoN 14.36 63.58 51.07 45.68
243 DMoN 13.84 63.82 56.68 49.40
16 DMoN 13.81 63.78 57.00 49.64
716 DMoN 13.91 63.70 55.59 49.16
383 DMoN 13.84 63.08 48.79 45.27
277 DMoN 14.06 63.70 46.72 39.63
274 DMoN 13.84 63.93 53.70 46.41
188 DMoN 13.94 63.70 55.66 49.31
796 DMoN 13.89 64.04 54.63 47.03
Mean 13.81 63.35 52.40 46.80
Variance 0.00 0.01 0.20 0.09
Table 38: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, and distance weight of 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (D) 16.08 60.64 49.75 44.05
550 DMoN–DPR (D) 15.42 56.02 50.59 52.21
243 DMoN–DPR (D) 15.78 58.21 58.58 53.21
16 DMoN–DPR (D) 15.91 54.69 53.21 54.59
716 DMoN–DPR (D) 15.85 57.35 50.14 49.28
383 DMoN–DPR (D) 18.82 55.05 51.96 48.46
277 DMoN–DPR (D) 19.89 54.78 50.10 47.08
274 DMoN–DPR (D) 15.03 56.06 55.73 54.53
188 DMoN–DPR (D) 13.71 59.75 61.71 54.99
796 DMoN–DPR (D) 18.02 56.74 58.17 51.22
Mean 16.45 56.93 53.99 50.96
Variance 0.04 0.04 0.18 0.13
Table 39: Results obtained by using DMoN–DPR pooling using the best variance weight of 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (V) 12.00 58.02 45.50 45.54
550 DMoN–DPR (V) 12.44 57.44 53.99 54.11
243 DMoN–DPR (V) 13.82 61.60 59.00 51.98
16 DMoN–DPR (V) 13.64 58.69 58.90 56.75
716 DMoN–DPR (V) 14.06 61.38 56.03 51.89
383 DMoN–DPR (V) 14.42 63.20 52.72 45.15
277 DMoN–DPR (V) 12.96 57.38 52.20 53.30
274 DMoN–DPR (V) 12.88 57.23 54.09 56.20
188 DMoN–DPR (V) 13.95 57.14 52.27 54.06
796 DMoN–DPR (V) 14.29 61.71 55.55 49.94
Mean 13.45 59.38 54.03 51.89
Variance 0.01 0.05 0.15 0.16
Table 40: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (E) 12.88 58.41 44.16 45.67
550 DMoN–DPR (E) 12.07 57.62 54.04 53.83
243 DMoN–DPR (E) 13.79 62.02 58.66 51.38
16 DMoN–DPR (E) 13.77 60.22 58.66 53.25
716 DMoN–DPR (E) 14.41 61.97 56.14 50.58
383 DMoN–DPR (E) 13.88 58.47 53.74 53.85
277 DMoN–DPR (E) 12.85 57.66 52.26 52.98
274 DMoN–DPR (E) 15.10 59.77 51.39 45.72
188 DMoN–DPR (E) 14.12 62.20 56.74 50.19
796 DMoN–DPR (E) 15.15 62.10 45.87 43.09
Mean 13.80 60.04 53.17 50.05
Variance 0.01 0.04 0.25 0.15
Table 41: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and variance weight 1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DV) 14.08 60.60 47.53 46.00
550 DMoN–DPR (DV) 14.07 55.94 54.18 57.79
243 DMoN–DPR (DV) 13.03 58.15 60.48 59.90
16 DMoN–DPR (DV) 12.03 55.51 54.29 57.77
716 DMoN–DPR (DV) 12.54 56.50 52.40 57.28
383 DMoN–DPR (DV) 13.82 56.10 52.59 57.34
277 DMoN–DPR (DV) 13.96 56.84 53.28 57.82
274 DMoN–DPR (DV) 12.40 56.41 52.47 57.72
188 DMoN–DPR (DV) 14.36 56.80 60.08 57.68
796 DMoN–DPR (DV) 14.34 55.61 53.64 54.26
Mean 13.46 56.85 54.09 56.36
Variance 0.01 0.02 0.14 0.15
Table 42: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DE) 15.49 55.62 48.15 53.66
550 DMoN–DPR (DE) 15.31 55.21 52.10 56.57
243 DMoN–DPR (DE) 13.74 56.71 59.84 58.07
16 DMoN–DPR (DE) 12.99 55.20 53.27 56.69
716 DMoN–DPR (DE) 14.00 56.38 59.34 58.29
383 DMoN–DPR (DE) 13.94 56.83 55.09 56.84
277 DMoN–DPR (DE) 14.97 54.26 54.14 60.85
274 DMoN–DPR (DE) 12.57 55.85 57.05 59.56
188 DMoN–DPR (DE) 12.99 56.54 58.60 58.86
796 DMoN–DPR (DE) 16.16 58.41 49.49 49.99
Mean 14.22 56.10 54.71 56.94
Variance 0.01 0.01 0.16 0.10
Table 43: Results obtained by using DMoN–DPR pooling using the best variance weight of 1, and entropy weight 0f 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (VE) 8.50 36.32 14.32 44.25
550 DMoN–DPR (VE) 13.12 57.28 52.98 56.87
243 DMoN–DPR (VE) 14.64 59.71 55.12 51.50
16 DMoN–DPR (VE) 13.92 58.45 62.37 58.80
716 DMoN–DPR (VE) 13.75 57.27 52.06 51.47
383 DMoN–DPR (VE) 14.04 57.77 56.72 58.23
277 DMoN–DPR (VE) 13.04 57.90 54.22 57.55
274 DMoN–DPR (VE) 12.10 56.56 51.63 56.60
188 DMoN–DPR (VE) 10.40 55.47 55.62 62.42
796 DMoN–DPR (VE) 15.01 60.44 46.37 47.48
Mean 12.85 55.72 50.14 54.52
Variance 0.04 0.49 1.75 0.32
Table 44: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, variance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
Seed Method Conductance Modularity NMI F1
993 DMoN–DPR (DVE) 15.43 55.58 47.68 53.96
550 DMoN–DPR (DVE) 14.27 55.01 53.99 58.87
243 DMoN–DPR (DVE) 12.76 56.10 53.71 58.34
16 DMoN–DPR (DVE) 14.62 53.36 49.11 57.45
716 DMoN–DPR (DVE) 13.06 54.46 57.73 62.18
383 DMoN–DPR (DVE) 12.92 55.12 58.39 60.97
277 DMoN–DPR (DVE) 14.04 55.08 51.68 58.26
274 DMoN–DPR (DVE) 12.59 55.68 58.34 60.52
188 DMoN–DPR (DVE) 11.41 57.40 63.84 62.83
796 DMoN–DPR (DVE) 13.51 54.94 48.67 55.26
Mean 13.46 55.27 54.31 58.86
Variance 0.01 0.01 0.27 0.08
Appendix C Code Implementation
The code was implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019), and was trained and evaluated using the evaluation protocols found in the official DMoN repository (Tsitsulin et al., 2023).

### v2

@@
-Deep Modularity Networks with Diversity–Preserving Regularization
+Deep Modularity Networks with Diversity-Preserving Regularization
 Abstract
 Graph clustering plays a crucial role in graph representation learning but often faces challenges in achieving feature-space diversity.
-While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they do not explicitly encourage diversity in the feature space among clusters.
-We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for intra-cluster diversity, and entropy-based for balanced assignments.
-Our method enhances clustering performance on benchmark datasets, namely Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics, achieving significant improvements in Normalized Mutual Information (NMI), and F1 scores.
-These results demonstrate the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters, especially in feature-rich datasets.
+While Deep Modularity Networks (DMoN) leverage modularity maximization and collapse regularization to ensure structural separation, they lack explicit mechanisms for feature-space separation, assignment dispersion, and assignment-confidence control.
+We address this limitation by proposing Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR), which introduces three novel regularization terms: distance-based for inter-cluster separation, variance-based for per-cluster assignment dispersion, and an assignment-entropy penalty with a small positive weight, encouraging more confident assignments gradually.
+Our method significantly enhances label-based clustering metrics on feature-rich benchmark datasets (paired two-tailed t-test, $p\leq 0.05$), demonstrating the effectiveness of incorporating diversity-preserving regularizations in creating meaningful and interpretable clusters.
 1 Introduction
-Graph clustering is a crucial problem within the graph representation learning domain, and is essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021).
+Graph clustering is a crucial problem within the graph representation learning domain, and essential for various applications, including but not limited to community detection in social networks (Perozzi et al., 2014; Xiao et al., 2015), data exploration (Perozzi and Akoglu, 2018), and functional module identification in biological networks (Jin et al., 2021).
 Consequently, there has been a surge in methods aimed at enhancing graph clustering performance.
@@
 In this context, graph pooling methods have also become prominent, as they provide a way to coarsen graphs by aggregating nodes into clusters (Cangea et al., 2018).
-Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly, or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
+Yet, some of these methods such as DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2019) were found to be computationally costly and/or too rigid, leading to poor convergence (Tsitsulin et al., 2023).
 To address these limitations, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) were introduced, combining spectral modularity maximization (Newman, 2006) with collapse regularization (Tsitsulin et al., 2023) to avoid trivial clustering solutions.
-While DMoN improves structural separation, it does not explicitly enforce diversity among clusters in the feature space.
-This limitation can result in structurally distinct but feature-wise homogeneous clusters, reducing effectiveness in applications requiring diverse and meaningfully differentiated clusters.
+Although DMoN captures structural communities, two practical gaps remain.
+First, while the encoder may use node attributes, the objective itself is agnostic to feature‑space geometry: it contains no term that directly rewards inter‑cluster feature separation or specialization of cluster assignments across nodes.
+This can yield clusters that are structurally distinct yet not well separated in feature space.
+Second, the objective lacks an explicit control for assignment confidence (e.g., entropy/temperature), so soft assignments can harden early.
+Why diversity matters.
+Accounting for both inter-cluster separation and intra-cluster variety prevents information collapse and yields richer representations.
+In practice, diversity improves downstream utility—for example, diversified recommendation in e-commerce optimizes cluster-level variety to boost engagement (Kim and Kang, 2025), and drawing items from distinct clusters raises serendipity without sacrificing accuracy (Aytekin and Karakaya, 2014).
+Beyond recommendation, biological networks exhibit the same principle: modules that are cohesive inside yet distinct outside reveal functional relationships that density-only objectives can miss (Spirin and Mirny, 2003).
+Recent graph-pooling work therefore introduces explicit mechanisms to spread centroids and diversify within-cluster features (Liu et al., 2023).
+These considerations motivate our design.
 Present Work
-To overcome the limitations of DMoN and enhance feature-space diversity in graph clustering, we propose an extension called Deep Modularity Networks with Diversity-Preserving Regularization (DMoN–DPR).
-Our method augments the original DMoN optimization objective by introducing additional regularization terms specifically designed to promote diversity among clusters.
-These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term, which maximizes the variance within each cluster’s assignment probabilities to capture a broader range of node features; and an entropy-based regularization term, which ensures a more balanced and uniform node distribution by maximizing the entropy of cluster assignments.
-Our proposed DMoN-DPR method addresses the shortcomings of existing approaches by explicitly enforcing feature-space diversity and ensuring that clusters remain well-separated, all while preventing them from becoming overly homogeneous in their node features.
-As a result, the clusters produced are both structurally cohesive and feature-wise distinct.
-Following the evaluation framework in (Tsitsulin et al., 2023), we demonstrate through extensive experiments that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
+Motivated by the above, we propose Deep Modularity Networks with Diversity-Preserving Regularization (DMoN-DPR), an extension of DMoN that augments its objective with three regularizers explicitly promoting diversity among clusters.
+These include a distance-based regularization term, which penalizes clusters with centroids too close in feature space to encourage distinct separation; a variance-based regularization term that increases the dispersion of assignment probabilities across nodes for each cluster, and an assignment‑entropy penalty added with a small positive weight, to avoid premature hard assignments and to let distance/variance terms drive diversity.
+By performing extensive evaluations, we demonstrate that our method improves clustering performance on the benchmark datasets, noticeably on Coauthor CS and Coauthor Physics datasets, which benefit from this enriched representation due to their rich feature spaces.
 2 Related Work
-Early methods for clustering on graphs often treated node features independently from graph structure.
-A straightforward example is k-means(features) (Lloyd, 1982), which clusters nodes solely based on attributes and disregards their connectivity.
-While simple, this approach often overlooks crucial structural relationships.
-To incorporate at least some graph information, k-means(DeepWalk) (Perozzi et al., 2014) combines node attributes with random-walk-based embeddings.
-This introduces a basic integration of graph structure but still relies on a two-step process of embedding and clustering.
-A more advanced alternative is k-means(DGI) (Tsitsulin et al., 2023), which employs node representations learned via Deep Graph Infomax (Veličković et al., 2018) and thereby captures both local and global context.
-However, because these k-means strategies only merge attributes and structure post hoc—rather than jointly optimizing them—they may lag behind end-to-end trainable methods that unify both graph topology and node features in a single framework.
-Moving toward more integrated approaches, Defferrard et al.
-(2016) introduced one of the first spectral graph convolution methods based on Chebyshev polynomials.
-Their technique approximates the spectral filters necessary for convolutional operations on graphs, reducing the computational cost of graph Fourier transforms and providing a scalable foundation for subsequent GNN designs.
-This seminal work illustrated how learned transformations on graph-structured data could drastically improve learning efficiency and accuracy, motivating the community to delve deeper into neural approaches.
-Beyond simply learning node-level representations, pooling mechanisms aim to capture hierarchical structures within graphs.
-Neural Overlapping Community Detection (NOCD) (Shchur and Günnemann, 2019) applies a probabilistic model that directly optimizes the likelihood of reconstructing the graph.
-Although effective in certain scenarios, NOCD may face scalability challenges and underutilizes node features in complex datasets.
-One of the earliest neural pooling frameworks, DiffPool (Ying et al., 2018), introduced a soft-clustering strategy that learns cluster assignments in an end-to-end manner.
-Despite its strong performance, its quadratic time complexity renders it less practical for large-scale graphs (Tsitsulin et al., 2023).
-To address computational bottlenecks and overcome potential overfitting, MinCut Pooling (Bianchi et al., 2020) reimagines the clustering problem through a spectral lens, incorporating normalized minCUT objectives with orthogonality constraints.
-While these constraints can improve cluster quality, they frequently lead to convergence issues.
-A related simplified variant, Ortho (Bianchi et al., 2020), focuses exclusively on orthogonality regularization, thus discarding many graph-specific objectives.
-This omission can produce suboptimal results, as the method forgoes some of the structural cues integral to good cluster formation.
-In another line of work, SAGPool (Lee et al., 2019) introduces a self-attention mechanism that ranks node importance, thereby selecting the most informative nodes during pooling.
-Although SAGPool effectively combines feature and topology considerations, computing attention coefficients can increase training costs, particularly for large graphs.
-Orthogonal to pooling-based approaches, Deep Graph Infomax (DGI) (Velickovic et al., 2019) pioneered an approach that maximizes mutual information between local substructures and a global summary of the graph.
-This method facilitates robust, label-agnostic node embeddings.
-The framework was later extended by InfoGraph (Sun et al., 2019), which applies a similar principle at the graph-level, yielding state-of-the-art results for unsupervised and semi-supervised tasks by capturing interactions between subgraphs and a holistic representation.
-Most recently, Deep Modularity Networks (DMoN) (Tsitsulin et al., 2023) have demonstrated a powerful alternative by directly incorporating modularity optimization with collapse regularization in an end-to-end trainable pipeline.
-DMoN bridges traditional graph community detection and neural feature extraction, achieving strong empirical performance on both synthetic and real-world graphs.
-By jointly optimizing for high modularity and balancing cluster assignment, DMoN outperforms many existing approaches in terms of normalized mutual information (NMI) and modularity scores, thus setting a new standard for clustering-oriented GNNs.
-3 Deep Modularity Networks with Diversity–Preserving Regularization
+Early graph clustering often decoupled features from structure.
+k-means on raw features (Lloyd, 1982) ignores connectivity; pairing k-means with DeepWalk or DGI embeddings (Perozzi et al., 2014; Veličković et al., 2018; Tsitsulin et al., 2023) injects structural signals but remains a two-stage pipeline, typically trailing end-to-end models that jointly learn representations and cluster assignments.
+Meanwhile, Chebyshev-based spectral convolutions (Defferrard et al., 2016) efficiently approximate graph filters, laying a scalable foundation for modern GNNs.
+On the other hand, pooling methods capture hierarchy.
+NOCD (Shchur and Günnemann, 2019) directly optimizes graph likelihood but can struggle with scale and feature use.
+DiffPool (Ying et al., 2018) learns soft cluster assignments end-to-end yet incurs quadratic cost (Tsitsulin et al., 2023).
+MinCut pooling (Bianchi et al., 2020) adds normalized cut and orthogonality terms that may hinder convergence; the Ortho variant keeps only the latter and loses structural cues.
+SAGPool (Lee et al., 2019) ranks node importance via self-attention, improving selectivity at added training cost on large graphs.
+For unsupervised representation learning, DGI (Velickovic et al., 2019) maximizes mutual information between local and global summaries; InfoGraph extends this idea to graph-level representations (Sun et al., 2019).
+More recently, DMoN (Tsitsulin et al., 2023) integrates modularity maximization with collapse regularization in an end-to-end framework, unifying community detection and neural feature learning and achieving strong NMI and modularity across benchmarks.
+3 Deep Modularity Networks with Diversity-Preserving Regularization
 DMoN tackles the issues of previous techniques by leveraging an optimization objective that combines insights from spectral modularity maximization (Newman, 2006) with a unique regularization term called collapse regularization (Tsitsulin et al., 2023).
@@
 The objective function is formulated as:
-$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
-where $\frac{1}{2m}\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, $\sqrt{\frac{k}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters.
+$L_{\text{DMoN}}(C;A)=\underbrace{-\frac{1}{2m}\text{Tr}(C^{\top}BC)}_{\text{Modularity term}}+\underbrace{{\frac{\sqrt{k}}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1}_{\text{Collapse Regularization term}}$ (2)
+where $(\frac{1}{2m})\text{Tr}(C^{\top}BC)$ is the modularity term (measuring how well clusters are internally connected compared to random chance), $B=A-\frac{dd^{\top}}{2m}$ is the modularity matrix, $d$ is the degree vector, $m$ is the total number of edges, ${\frac{\sqrt{k}}{n}}\left\|\sum_{i}C_{i}^{\top}\right\|_{F}-1$ represents the collapse regularization term, which discourages the formation of trivial clusters by penalizing the Frobenius norm of the assignment matrix $C$, $\|\cdot\|_{F}$ denotes the Frobenius norm, and $k$ is the number of clusters.
 This regularization term encourages balanced cluster assignments, thereby improving the quality of clustering and avoiding degenerate solutions.
-Building upon the DMoN framework, we augment the DMoN objective function $L_{\text{DMoN}}$ with three new regularization terms, resulting in the following enhanced objective function:
-$L_{\text{DMoN--DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\cdot L_{\text{DPR}}^{\text{Distance}}+W_{\text{var}}\cdot L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\cdot L_{\text{DPR}}^{\text{entropy}}$ (3)
-Here, $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$ are hyperparameters that control the influence of each regularization term.
-The additional terms are designed to promote inter-cluster separation, intra-cluster diversity, and balanced cluster assignments.
+Building upon the DMoN framework, we augment the objective with three regularizers:
+$L_{\text{DMoN-DPR}}(C;A)=L_{\text{DMoN}}+W_{\text{dist}}\,L_{\text{DPR}}^{\text{distance}}+W_{\text{var}}\,L_{\text{DPR}}^{\text{variance}}+W_{\text{entropy}}\,L_{\text{DPR}}^{\text{entropy}}.$ (3)
+Here, $W_{\text{dist}}$, $W_{\text{var}}$, and $W_{\text{entropy}}$ control the influence of each term.
+These additions promote inter-cluster separation (distance) and per-cluster assignment dispersion (variance).
+We use a small positive entropy weight so per-node entropy decreases slowly, preserving higher uncertainty early and thus indirectly encouraging more balanced cluster usage during exploration; the primary mechanism for cluster-size balance remains DMoN’s collapse regularizer.
 Distance-Based Regularization
@@
 $L_{\text{DPR}}^{\text{Distance}}=\frac{1}{k(k-1)}\sum_{i=1}^{k}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{k}\text{ReLU}\left(\epsilon-\|\mu_{i}-\mu_{j}\|_{2}^{2}\right)$ (4)
-where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss.
+where $k$ is the number of clusters, $\mu_{i}$ and $\mu_{j}$ are the centroids of clusters $i$ and $j$, computed as $\mu_{i}=\frac{\sum_{v=1}^{n}C_{vi}X_{v}}{\sum_{v=1}^{n}C_{vi}}$, where $X_{v}$ is the feature vector of node $v$, $\|\cdot\|_{2}^{2}$ denotes the squared Euclidean norm, $\epsilon$ is a predefined threshold that sets the minimum acceptable squared distance between cluster centroids, and ReLU is the Rectified Linear Unit function, ensuring that only distances below $\epsilon$ contribute to the loss.
 By penalizing pairs of clusters whose centroids are closer than $\epsilon$, clusters can be effectively pushed apart in the feature space.
@@
 Variance-Based Regularization
-The variance-based regularization term aims to increase the diversity within each cluster by maximizing the variance of node assignments in each cluster:
+The variance-based regularization term encourages dispersion in the assignment matrix by maximizing, for each cluster, the variance of its assignment probabilities across nodes:
 $L_{\text{DPR}}^{\text{variance}}=-\frac{1}{k}\sum_{i=1}^{k}\text{Var}(C_{:i})$ (5)
-[…]
+where $\text{Var}(C_{:i})$ denotes the variance of the assignment probabilities of all nodes to cluster $i$.
+Maximizing this variance prevents uniform cluster columns (e.g., every node assigned with equal probability), ensuring that clusters specialize over different subsets of nodes and thereby enhancing assignment diversity.
 Entropy-Based Regularization
-The entropy-based regularization term promotes a uniform distribution of nodes across clusters, preventing any single cluster from dominating. It is formulated as:
+We define:
 $L_{\text{DPR}}^{\text{entropy}}=-\frac{1}{n}\sum_{v=1}^{n}\sum_{i=1}^{k}C_{vi}\log(C_{vi}+\delta)$ (6)
-where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability.
-By maximizing the entropy of the cluster assignments, this term encourages uncertainty in the assignments, promoting a more balanced distribution of nodes among clusters.
-This prevents scenarios where one cluster might contain a majority of the nodes, thus enhancing overall diversity and cluster balance.
-Therefore, by integrating these diversity-preserving regularization terms into the DMoN objective, we can simultaneously consider graph structure, cluster balance, inter-cluster separation, and intra-cluster diversity.
+which is the average per‑node Shannon entropy of the soft assignments, where $n$ is the number of nodes, $C_{vi}$ is the soft assignment probability of node $v$ to cluster $i$, $\delta$ is a small constant added to prevent logarithm of zero, ensuring numerical stability.
+We minimize $L_{\text{DPR}}$ to gently encourage more confident assignments.
+Because minimizing entropy can lead to premature hardening, we apply a small weight $W_{\text{entropy}}$ (typically 0.001–0.1) so that entropy decreases slowly during training.
 4 Experiments
-Datasets
-To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on the widely used Cora, CiteSeer, and PubMed datasets Yang et al.
-(2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018).
+Evaluation Protocol
+To evaluate the effectiveness of our proposed regularization objective, we conducted experiments on Cora, CiteSeer, and PubMed datasets (Yang et al., 2016), as well as the Coauthor CS and Coauthor Physics datasets (Shchur et al., 2018).
+Information regarding these datasets is presented in Appendix A.
+Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure.
+Additionally, we use the following quantitative measures to analyze feature‐space diversity: average inter‐centroid distance (mean pairwise Euclidean distance between cluster centroids), minimum inter‐centroid distance (the smallest pairwise distance among centroids), average intra‐cluster variance (average within‐cluster variance of node embeddings), and Silhouette score (a standard measure of per‐point cohesion vs.
+separation).
+The quantification of diversity results are presented in Appendix B.
+Baseline & Implementation Details
+We select DiffPool (Ying et al., 2018), MinCut pooling (Bianchi et al., 2019) and DMoN (Tsitsulin et al., 2023) as our baselines.
+Additionally, we focus on ablation studies and significance tests against vanilla DMoN to isolate the impact of the new diversity terms.
+Furthermore, similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by a pooling layer.
+Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001.
+The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– is available at www.github.com/YasminSalehi/DMoN-DPR.
+5 Results
+Table 1: Comparison of clustering methods on three datasets (Cora, CiteSeer, PubMed). Values are in percentage.
+Method Cora CiteSeer PubMed
+Graph Labels Graph Labels Graph Labels
+$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
+DiffPool 15.99 62.78 40.13 46.55 7.92 66.69 33.40 47.83
+MinCut 13.65 71.79 37.74 39.15 6.19 75.21 25.21 35.28 11.14 54.66 22.48 41.31
+DMoN 10.53 72.77 43.92 46.93 4.86 75.45 29.67 42.46 8.61 57.13 22.39 43.23
+DPR(D) 10.95 72.20 43.98 47.51 4.95 75.56 30.09 42.46 8.60 57.14 22.38 43.24
+DPR(V) 11.87 70.54 43.34 46.39 4.95 75.17 29.76 42.85 8.61 57.13 22.39 43.23
+DPR(E) 10.49 72.77 44.00 47.02 4.93 75.25 29.95 42.96 8.62 57.11 22.36 43.22
+DPR(DV) 10.89 72.07 44.40 47.36 5.09 75.18 30.50 43.25 8.60 57.14 22.38 43.24
+DPR(DE) 10.88 72.30 44.30 47.53 5.07 75.18 30.47 43.13 8.60 57.12 22.38 43.25
+DPR(VE) 11.86 70.51 43.24 46.36 5.05 75.01 30.17 43.27 8.60 57.13 22.37 43.21
+DPR(DVE) 10.93 72.05 44.37 47.32 5.16 75.00 30.25 43.00 8.61 57.11 22.36 43.24
+Table 2: Comparison of clustering methods on two datasets (Coauthor CS, Coauthor Physics). Values are in percentage.
+Method Coauthor CS Coauthor Physics
+Graph Labels Graph Labels
+$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
+DiffPool 18.19 62.24 53.92 53.47 13.91 57.44 56.22 51.71
+MinCut 21.36 71.58 64.33 49.00 13.93 61.75 51.39 47.79
+DMoN 18.63 72.60 69.26 59.26 13.70 63.45 53.50 47.51
+DPR(D) 18.89 72.30 71.17 61.82 16.44 56.89 53.49 50.99
+DPR(V) 18.75 72.50 69.56 59.94 13.18 58.59 53.93 52.78
+DPR(E) 19.85 70.92 71.58 61.33 13.49 59.33 52.83 51.09
+DPR(DV) 19.14 71.96 70.72 61.35 13.43 56.48 55.84 57.99
+DPR(DE) 19.81 70.57 70.96 62.36 14.17 56.89 54.02 55.02
+DPR(VE) 19.89 70.81 71.47 61.33 13.11 55.88 49.95 53.58
+DPR(DVE) 19.69 70.71 71.28 62.67 12.84 55.47 53.50 57.96
+Table 3: t-statistics and p-values for all datasets (rounded to two decimal places).
+Metric Cora CiteSeer PubMed Coauthor CS Coauthor Physics
+t-stat p-val t-stat p-val t-stat p-val t-stat p-val t-stat p-val
+Conductance -0.88 0.40 -1.62 0.14 0.37 0.72 -5.25 $1e^{-6}$ 0.51 0.62
+Modularity 1.81 0.10 1.17 0.27 0.21 0.84 5.88 $1e^{-6}$ 17.56 $1e^{-6}$
+NMI -0.63 0.54 -1.56 0.15 0.23 0.83 -4.64 $1e^{-6}$ -2.49 0.03
+F1 -0.40 0.70 -1.28 0.23 -0.95 0.37 -2.51 0.03 -14.65 $1e^{-6}$
+5.1 Effectiveness of DMoN-DPR in Clustering
+To assess the effectiveness of DMoN-DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 randomly selected seeds achieved with DMoN-DPR (denoted as DPR in all the tables) to those obtained with our baselines.
+Tables 1 (Cora, CiteSeer, PubMed) and 2 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
+For completeness, we defer supporting analyses to the appendix: (i) quantitative diagnostics of feature-space diversity (average/min inter-centroid distance, average intra-cluster variance, Silhouette score) (Appendix B), (ii) t–SNE visualizations of the learned clusters versus ground-truth labels (Appendix C), (iii) ablation and hyperparameter-selection studies that isolate the effect of each DPR term (Appendix D), and (iv) per-seed scores and full tables with error bars (Appendix E).
+Cora, CiteSeer, and PubMed.
+Overall, the results confirm that adding diversity-preserving regularizers to DMoN almost never harms the purely topological scores—conductance ($C$ ↓) and modularity ($Q$ ↑)—while consistently lifting label-aware metrics such as NMI and F1.
+On Cora, DPR(DV) achieves the best NMI (44.40%) and only narrowly trails DPR(DE) on F1.
+On CiteSeer, although DiffPool attains the highest NMI (33.40%) and F1 (47.83%), DPR variants remain competitive without the instability that causes DiffPool to collapse on PubMed  [^1] With a single DiffPool layer and our fixed cluster budget ($C=3$), the link-prediction loss on the 19 k-node PubMed graph stagnated and the assignment matrix collapsed to one cluster.
+Increasing the budget restores numerical stability, but yields only marginal label quality and would also violate our fixed-$C$ protocol, so we omit DiffPool scores for PubMed..
+Among the DMoN-family methods, DPR(DV) is the strongest on CiteSeer.
+On PubMed, all DPR variants remain stable, in contrast to DiffPool, while MinCut yields the highest NMI (22.48%) and DPR(DE) edges out others in F1.
+The fact that DPR(D) and DPR(DV) match the best conductance (8.60%) confirms that the regularizers maintain cut quality.
+Coauthor CS and Coauthor Physics.
+On the more feature-rich [^2] We define feature-rich graphs as those that exhibit (i) high feature dimensionality and (ii) high average Shannon entropy per node.
+Please see Appendix A for more detail.
+Coauthor graphs, diversity-preserving regularization brings substantial gains.
+On Coauthor CS, DPR(DVE) achieves a new best in F1 (62.67%) while DPR(E) sets the highest NMI (71.58%), both outperforming the already strong DMoN baseline without sacrificing conductance or modularity.
+On Coauthor Physics, DPR(DVE) obtains the lowest conductance (12.84%) and, along with DPR(DV), surpasses 57% F1—demonstrating that the full three-term objective generalizes well to larger, noisier graphs.
+Key Insights.
+Therefore, our results show that the effects of distance-based (D), variance-based (V), and entropy-based (E) regularizations depend on dataset feature richness.
+On feature-rich datasets like Coauthor CS and Physics, adding these terms—especially D—significantly boosts NMI and F1, with DMoN-DPR(DV) raising F1 by over 10 percentage points.
+The V term increases assignment dispersion across nodes and performs best when combined with D.
+The E term, used with a small positive weight, gradually sharpens per-node assignments, which can indirectly improve cluster balance during early exploration on feature-rich datasets, but offers little benefit on simpler datasets like PubMed; the collapse regularizer remains the primary mechanism for cluster-size balance.
+Overall, combining D and V often yields the highest performance, underscoring the need to tailor regularization to dataset characteristics for optimal clustering.
+Statistical Significance.
+To strengthen our claims, we conduct a paired t‐test on Conductance, Modularity, NMI, and F1, comparing DMoN vs.
+the best‐performing DMoN–DPR variants.
+We ensure both methods use exactly the same seeds in each dataset.
+The results are summarized in Table 3.
+On the citation benchmarks we adopt the strongest variant for each graph (DV on Cora and CiteSeer, DVE on PubMed).
+Although the average NMI and F1 of DMoN–DPR surpass or match the vanilla DMoN baseline, the paired two-tailed $t$-tests in Table 3 show p-values above 0.10.
+This indicates that, given the modest sample of random seeds, the improvements are encouraging but not yet conclusive for these sparsely featured graphs.
+The picture is markedly different on the feature-rich coauthor networks.
+The gains on NMI and F1 metrics are highly significant, indicated by $p\leq 0.05$, as seen in Table 3.
+These results validate our hypothesis that encouraging latent cluster dispersion is most beneficial when node attributes are abundant and heterogeneous, and they confirm that DMoN–DPR maintains the stability advantage of the original objective while translating it into measurably better community recovery.
+Trade-Offs Between Structural and Label-Based Metrics.
+Our results reveal a clear trade-off between structural metrics (e.g., Modularity, Conductance) and label-based metrics (e.g., NMI, F1).
+Diversity-preserving regularization improves label alignment, especially on feature-rich datasets like Coauthor CS and Physics, by forming clusters that better match ground-truth labels—though sometimes at the cost of structural cohesion (e.g., lower modularity).
+Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest.
+As a result, DMoN-DPR emerges as a more effective choice for supervised tasks prioritizing NMI and F1, whereas vanilla DMoN remains competitive for unsupervised scenarios where conductance and modularity are paramount.
+Tailoring the clustering strategy to the dataset and task requirements is therefore essential.
+5.2 Execution Times and Runtime Analysis
+Table4 reports the execution times (in seconds) of DMoN and DMoN–DPR on five commonly used graph datasets.
+All experiments were conducted on a CPU to ensure consistent runtime measurement without GPU scheduling effects.
+Table 4: Execution times (in seconds) on CPU for DMoN and DMoN–DPR.
+Method Cora CiteSeer PubMed Coauthor CS Coauthor Physics
+DMoN 32 76 310 722 2107
+DMoN–DPR 35 71 300 751 1946
+Although DMoN–DPR includes additional regularization terms for Distance-based, Entropy-based, and Variance-based constraints, its runtime is sometimes comparable to or even marginally lower than plain DMoN.
+The reason is that the most time-consuming part of both methods is dominated by adjacency-based operations (e.g., matrix multiplications with the graph Laplacian or adjacency matrix), which scale on the order of $\mathcal{O}(|E|\,k)$ or $\mathcal{O}(N^{2}k)$ for large, dense graphs (where $N$ is the number of nodes, $E$ is the set of edges, and $k$ is the number of clusters).
+In contrast, the extra DPR calculations—computing entropy across nodes, the variance of cluster assignments, or centroid distances—represent only $\mathcal{O}(N\,k+k^{2}F)$ overhead (with $F$ being the feature dimension) and are typically negligible next to the larger matrix multiplications.
+5.3 Limitations and Future Directions
+While DMoN-DPR introduces clear benefits, it also presents a few limitations.
+First, tuning the diversity-preserving weights for distance, variance, and entropy remains manual.
+Though our empirical studies revealed consistent patterns: high-dimensional datasets like Coauthor CS and Physics benefited from larger distance weights (1 or 10), variance weights of 1 or 0.1 worked well across most datasets, and entropy value of 0.1 served as a good upper bound.
+A promising direction for future work is to develop an adaptive scheme that learns these weights during training, reducing reliance on manual tuning.
+Second, the gains from diversity regularization are more pronounced on feature-rich datasets.
+On lower-dimensional datasets like PubMed, improvements in NMI and F1 are modest—though visualizations (Figure 1) suggest that DMoN–DPR still forms more coherent clusters than the baseline.
+Future research could explore alternative regularizers to better capture structure in such settings.
+6 Conclusion
+In this work, we presented DMoN–DPR, an enhanced version of Deep Modularity Networks (DMoN) that incorporates diversity-preserving regularizations to enrich the clustering objective.
+By introducing distance-, entropy-, and variance-based penalties alongside the original modularity and collapse regularization terms, our approach promotes diversity by increasing inter-cluster feature separation (distance) and within-cluster assignment dispersion (variance), while a small-weight entropy term gradually sharpens assignments without sacrificing exploration.
+Our empirical results on the Coauthor CS and Coauthor Physics benchmark datasets indicate that DMoN–DPR significantly improves alignment with ground-truth labels in feature-rich datasets.
+While structural metrics such as modularity and conductance remained competitive, the additional diversity constraints contributed to more interpretable and semantically meaningful cluster formations, marked by significant improvements in NMI and F1 scores, producing higher alignment with ground‐truth labels and generally reflecting clearer semantic distinctions.
+Overall, DMoN-DPR achieves a balance between interpretability and structural integrity, making it a powerful solution for graph clustering tasks that demand both topological cohesion and semantic differentiation, marking a significant advancement in feature-aware graph clustering.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We identified a key limitation of DMoN—namely, that it often yields clusters with insufficient feature‐space diversity—and addressed it by introducing three novel regularization terms that are distance, variance, and entropy based.
+In our paper, we have shown that the impact on label‐based metrics can be substantial on feature‐rich datasets such as Coauthor CS and Coauthor Physics, where DMoN–DPR significantly improves NMI and F1 compared to baseline DMoN.
+This demonstrates that our approach works especially well in scenarios where node attributes carry high signal, thus expanding DMoN’s applicability to more attribute‐driven clustering tasks.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer:[Yes]
+  9.
+Justification: In Section 5.1, we clearly highlight the trade-off between DMoN and DMoN–DPR, outlining the scenarios in which each is most suitable.
+Furthermore, Section 5.2 provides a detailed discussion of the limitations of our approach and potential directions for future work.
+We have also included a runtime analysis in Appendix D.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14.
+Justification: Our main contribution is an empirical extension of DMoN that aims to preserve feature diversity, rather than a new theoretical framework.
+Tsitsulin et al.
+(2023) showed that DMoN’s collapse regularization avoids trivial solutions and is asymptotically consistent under certain generative assumptions.
+Our distance-, variance-, and entropy-based regularizers further promote diverse, well-separated clusters without disrupting DMoN’s convergence.
+Each term is differentiable (or piecewise differentiable), ensuring compatibility with gradient-based optimization.
+Specifically:
+    • $R_{D}$ (distance) repels centroids in feature space, reducing cluster overlap.
+    • $R_{V}$ (variance) prevents overly tight clusters, mitigating premature “hard” assignments.
+    • $R_{E}$ (entropy) avoids early dominance by any single cluster, helping to avert mode collapse.
+    These terms add complementary constraints but do not create new problematic local minima or break DMoN’s consistency.
+Rather, they refine the solution space by adding complementary constraints.
+While we do not offer formal proofs here, we consistently observe stable training across all studied datasets.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: All ablation studies, seed values, and hyperparameter settings are thoroughly documented in the appendix.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The datasets used in this paper are standard benchmarks, all publicly available through the PyTorch Geometric library. The code will also be released publicly.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All the details have been discussed in the appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: The result for all trial runs as well as their mean and standard deviation has been reported in the appendix.
+Additionally, statistical significance testing has also been reported in the Results (Section 5.1).
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The hardware and implementation settings have been mentioned in the appendix as well as in Section 4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The work uses public, consented academic-citation and co-authorship data and involves no sensitive attributes or human subjects.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49.
+Justification: Improved graph clustering can aid knowledge discovery in scientific networks (positive), but could also facilitate surveillance or profiling if applied to social graphs (negative).
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The released code trains only on publicly available graphs and does not include any pre-trained model with dual-use risk.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The paper cites PyTorch Geometric and references each dataset with its original publication and Creative Commons/licensing where applicable.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: No new datasets or pre-trained models are released; only source code is provided.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The study uses only machine-generated data; no human participants were involved.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Not applicable—no human-subject research.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: No large-language models are part of the methodology; any text editing assistance was purely editorial.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A Datasets
 The Cora, CiteSeer, and PubMed datasets are citation networks where nodes represent papers and edges denote citation relationships, with node labels corresponding to the topic of each paper.
 The Coauthor CS and Coauthor Physics datasets are co-authorship graphs, where nodes represent authors and edges indicate collaborations, with labels reflecting the field of research.
-The key characteristics of these datasets are summarized in Table 1.
+The following table provides information about the datasets used for running experiments.
+Feature-rich Datasets
+A graph is considered to be feature-rich when its nodes are described by high-dimensional feature vectors that carry substantial information content.
+Formally, for each node $v$, the feature vector $x_{v}$ is normalized into a probability distribution $p_{v}=\frac{x_{v}}{\mathbf{1}^{\top}x_{v}},$ and its information content is quantified by the Shannon entropy $H(p_{v})=-\sum_{i=1}^{F}p_{v,i}\log_{2}p_{v,i}.$ A dataset of such nodes is said to be feature-rich when it combines two key properties: (1) high dimensionality, meaning that the number of features $|X|=F$ is large, and (2) high average entropy, defined as:
+$\frac{1}{N}\sum_{v}H(p_{v}),$
+indicating that on average, the features across nodes are diverse and well-distributed rather than concentrated in a few dimensions.
+Together, these conditions ensure that the feature space provides both breadth and depth, enabling richer representations for learning tasks.
+For example, Coauthor‑CS is a feature-rich dataset for having $|X|=6805$ and entropy = 5.51 bits, versus PubMed with $|X|=500$ and entropy = 5.17 bits.
 Table 5: Dataset characteristics, where ($|V|$) is the number of vertices , ($|E|$) is the number of edges, $|X|$ is the number of features, and $|Y|$ is the number of cluster labels.
-Dataset $|V|$ $|E|$ $|X|$ $|Y|$
-Cora 2708 5278 1433 7
-CiteSeer 3327 4614 3703 6
-PubMed 19717 44325 500 3
-Coauthor CS 18333 81894 6805 15
-Coauthor Phys 34493 247962 8415 5
-Evaluation Protocol
-Following the evaluation protocol outlined in (Tsitsulin et al., 2023), we employed the following metrics to assess performance: graph conductance (C), modularity (Q), Normalized Mutual Information (NMI) with ground-truth labels, and the pairwise F1 measure.
-These metrics collectively capture the structural quality of clusters and their alignment with ground-truth labels.
-Baseline
-We select DMoN as our sole baseline because it consistently achieves state-of-the-art results across multiple datasets and metrics, as shown in prior work (Tsitsulin et al., 2023).
-It outperforms a wide range of established methods—such as k-means(features) (Lloyd, 1982), k-means(DeepWalk) (Perozzi et al., 2014), k-means(DGI) (Veličković et al., 2018), NOCD (Shchur and Günnemann, 2019), DiffPool (Ying et al., 2018), MinCut Pooling (Bianchi et al., 2020), and Ortho (Bianchi et al., 2020)—each representing distinct approaches that prioritize node attributes, graph structure, or both.
-Unlike traditional clustering methods like k-means (which lacks graph structure awareness) or embedding-based techniques (which emphasize representation learning over explicit clustering), DMoN offers an end-to-end differentiable framework that directly optimizes modularity and uses collapse regularization to avoid trivial solutions.
-Furthermore, it exhibits superior scalability and computational efficiency compared to pooling-based approaches such as DiffPool and MinCutPool, making it particularly suited for the large, feature-rich datasets in our experiments.
-By focusing on feature-space diversity within DMoN’s modularity-driven clustering paradigm, we ensure that our enhancements remain aligned with state-of-the-art structural considerations, providing a clear and meaningful evaluation of the improvements introduced by our proposed modifications.
-Implementation Details
-Similar to (Tsitsulin et al., 2023), one layer of GCN (Kipf and Welling, 2016) with 512 neurons was used to create the graph embedding, followed by the DMoN/DMoN–DPR pooling layer, as illustrated in Figure 1.
-Likewise, the models were trained for 1000 epochs using Adam optimizer with a learning rate of 0.001.
-The experiments were run on an Apple M2 Max CPU.
-The code–implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019)– will be made publicly available at www.github.com/TBA.
-Figure 1: Overview of the DMoN/DMoN–DPR pipeline.
-Hyperparameter Tuning
+Dataset $|V|$ $|E|$ $|X|$ $|Y|$ Mean Entropy (bits)
+Cora 2708 5278 1433 7 4.05
+CiteSeer 3327 4614 3703 6 4.94
+PubMed 19717 44325 500 3 5.17
+Coauthor CS 18333 81894 6805 15 5.51
+Coauthor Physics 34493 247962 8415 5 4.76
+Appendix B Quantification of Diversity
+Table 6: Comparison of diversity resulted from the clustering methods on three datasets (Cora, CiteSeer, PubMed).
+AICD is average inter-centroid distance, MICD is minimum inter-centroid distance, AICV is average intra-cluster variance, and Sil is the Silhouette score.
+Method Cora CiteSeer PubMed
+AICD MICD AICV Sil AICD MICD AICV Sil AICD MICD AICV Sil
+DiffPool 3.69 2.62 0.02 0.16 5.03 4.13 0.02 0.25
+MinCut 6.37 5.42 0.02 0.36 9.03 8.18 0.03 0.46 3.03 2.62 0.00 0.46
+DMoN 8.10 5.14 0.04 0.27 10.11 7.34 0.04 0.37 2.77 2.59 0.00 0.48
+DPR(D) 8.01 4.99 0.04 0.26 10.15 7.33 0.04 0.37 2.77 2.59 0.00 0.48
+DPR(V) 8.93 5.25 0.05 0.23 10.20 7.17 0.05 0.35 2.77 2.59 0.00 0.48
+DPR(E) 8.12 5.14 0.04 0.2 10.31 7.14 0.05 0.35 2.77 2.59 0.00 0.48
+DPR(DV) 8.12 4.80 0.04 0.26 10.26 7.26 0.05 0.35 2.77 2.59 0.00 0.48
+DPR(DE) 8.03 4.88 0.04 0.26 10.35 7.23 0.05 0.35 2.78 2.60 0.00 0.48
+DPR(VE) 8.93 5.25 0.05 0.23 10.39 7.12 0.05 0.34 2.77 2.59 0.00 0.48
+DPR(DVE) 8.13 4.81 0.04 0.26 10.43 7.07 0.05 0.33 2.78 2.60 0.00 0.48
+Table 7: Comparison of diversity resulted from the clustering methods on two datasets (Coauthor CS, Coauthor Physics).
+AICD is average inter-centroid distance, MICD is minimum inter-centroid distance, AICV is average intra-cluster variance, and Sil is the Silhouette score.
+Method Coauthor CS Coauthor Physics
+AICD MICD AICV Sil AICD MICD AICV Sil
+DiffPool 6.05 2.93 0.03 0.12 4.35 2.14 0.01 0.23
+MinCut 10.46 6.89 0.06 0.27 10.88 8.56 0.04 0.42
+DMoN 20.47 8.15 0.26 0.17 13.97 10.51 0.09 0.37
+DPR(D) 17.72 7.01 0.19 0.18 14.55 7.27 0.11 0.23
+DPR(V) 21.02 8.33 0.27 0.17 17.44 9.79 0.16 0.27
+DPR(E) 31.67 11.14 0.58 0.15 14.72 7.46 0.11 0.25
+DPR(DV) 18.16 6.94 0.20 0.17 14.72 7.46 0.11 0.25
+DPR(DE) 24.48 9.38 0.37 0.19 14.58 7.66 0.11 0.25
+DPR(VE) 31.75 11.17 0.59 0.15 19.16 11.15 0.19 0.29
+DPR(DVE) 24.66 9.79 0.36 0.18 14.31 7.73 0.11 0.26
+The diversity diagnostics in Tables 67 echo the accuracy gains: augmenting DMoN with our diversity-preserving regularizers widens the geometric spread of clusters—captured by the average inter-centroid distance (AICD) and minimum inter-centroid distance (MICD)—while largely preserving internal cohesion, as indicated by the near-constant average intra-cluster variance (AICV) and Silhouette score (Sil).
+On the citation graphs, most DPR variants leave AICD and MICD within a few percent of the DMoN baseline (the largest jump is $\approx 10\%$ on Cora for V/VE), while AICV and Silhouette score remain virtually unchanged, confirming that clusters are not overstretched when the feature signal is weak.
+The feature-rich Coauthor graphs display a stronger effect: the entropy-weighted VE model, for example, boosts AICD on Coauthor CS from 20.5 to 31.8 and MICD from 8.2 to 11.2—at the cost of higher AICV—while hybrids that include the distance term (DE, DVE) achieve a better trade-off, enlarging centroid spacing by $15-25\%$ but capping the rise in AICV so that Silhouette score stays on par with, or slightly above, the DMoN baseline.
+The sole exception is Coauthor Physics, where Silhouette score dips because DPR variants both push centroids outward and allow somewhat broader clusters; this reduces compactness even as separation improves.
+Overall, DPR regularization fulfills its design goal: clusters become globally more dispersed yet remain locally cohesive, yielding modest benefits on low-feature graphs and pronounced gains when node attributes are rich and heterogeneous.
+Note that AICV measures embedding variance within clusters, whereas our V term maximizes assignment variance across nodes; any changes in AICV under V are therefore indirect.
+Appendix C Visualization of Clusters
+Figure 1: Visualization of clusters (circles) vs.
+ground truth labels (crosses) formed by DMoN on (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics, and by DMoN–DPR on (f) Cora, (g) CiteSeer, (h) PubMed, (i) Coauthor CS, and (j) Coauthor Physics datasets.
+The clustering results visualized in Figure 1, obtained using t–SNE with a fixed perplexity (30) and learning rate (200), demonstrate that DMoN–DPR generally outperforms DMoN across the benchmark datasets Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics.
+For Cora and CiteSeer datasets, DMoN–DPR produces tighter, more well-separated clusters with improved alignment to the ground truth labels, contrasting with the more scattered and less defined clusters formed by DMoN.
+On the other hand, PubMed remains challenging for both methods, resulting in notable overlap and scattered clusters; yet, DMoN–DPR still provides slightly tighter groupings.
+Turning to the Coauthor CS and Coauthor Physics datasets, a similar trend emerges.
+While DMoN adequately captures overarching modular structures, it can occasionally struggle to separate nuanced subgroups.
+In contrast, DMoN–DPR consistently yields more cohesive and better-separated clusters, offering clearer decision boundaries and higher alignment with the true labels.
+This improvement is particularly notable in Coauthor CS, where DMoN–DPR visibly reduces overlap and sharpens cluster delineations.
+These findings underscore the effectiveness of promoting diversity within hidden representations, as DMoN–DPR not only excels in forming compact and accurate clusters for datasets with greater feature variability but also remains competitive on more challenging datasets.
+Appendix D Ablation Study
+D.1 Hyperparameter Tuning
+DiffPool and MinCut
+For DiffPool, the entropy weight was set to $1\times 10^{-5}$ on the Coauthor Physics dataset to achieve optimal performance, and to $1\times 10^{-4}$ for all other datasets.
+For MinCut pooling, both the mincut loss and orthogonality loss weights were set to 1, which consistently yielded the best results.
+DMoN–DPR
 The weighting coefficients $W_{\text{dist}}$, $W_{\text{entropy}}$, and $W_{\text{var}}$, as well as $\epsilon$ are hyperparameters that need to be tuned based on the specific dataset and desired clustering behavior.
 They control the trade-off between structural modularity and diversity preservation.
-To optimize these hyperparameters, we conducted evaluations on the Cora, CiteSeer, PubMed, Coauthor CS and Coauthor Physics datasets using 10 random seeds, similar to how it was done in (Tsitsulin et al., 2023).
+To optimize these hyperparameters, we conducted evaluations on the Cora, CiteSeer, PubMed, Coauthor CS and Coauthor Physics datasets using 10 random seeds, similar to how it was done in (Tsitsulin et al., 2023), selected using a random number generator.
 The seeds that resulted in the best performance when using the DMoN pooling layer were selected.
@@
 This systematic approach allowed us to assess the impact of each regularization term and their combinations on model performance.
-Ablation studies are presented in Section A.
-5 Results
-Table 2: Comparison of clustering methods on three datasets (Cora, Citeseer, Pubmed). Values are in percentage.
-Method Cora Citeseer Pubmed
-Graph Labels Graph Labels Graph Labels
-$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
-DMoN 11.00 72.39 43.65 46.24 4.92 75.72 30.26 42.17 8.66 57.03 22.68 45.18
-DMoN–DPR(D) 10.23 72.97 45.31 50.00 4.75 75.64 30.61 42.76 8.66 57.02 22.68 45.19
-DMoN–DPR(V) 11.64 70.80 44.27 48.18 4.93 75.59 30.46 42.26 8.66 57.02 22.68 45.18
-DMoN–DPR(E) 11.00 72.37 43.63 46.25 5.02 75.43 30.52 42.54 8.67 57.02 22.68 45.19
-DMoN–DPR(DV) 10.37 72.68 45.41 50.13 4.76 75.51 30.59 42.85 8.66 57.02 22.68 45.19
-DMoN–DPR(DE) 10.33 72.88 45.18 49.88 4.86 75.35 30.77 43.27 8.66 57.02 22.67 45.19
-DMoN–DPR(VE) 11.64 70.77 44.15 48.09 5.09 75.26 30.86 42.86 8.66 57.01 22.68 45.18
-DMoN–DPR(DVE) 10.41 72.65 45.43 50.06 4.85 75.14 30.77 43.42 8.66 57.02 22.67 45.18
-Table 3: Comparison of clustering methods on two datasets (Coauthor CS, Coauthor Physics). Values are in percentage.
-Method Coauthor CS Coauthor Phys
-Graph Labels Graph Labels
-$C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$ $C\downarrow$ $Q\uparrow$ NMI $\uparrow$ F1 $\uparrow$
-DMoN 18.80 72.43 69.45 58.83 13.81 63.35 52.40 46.80
-DMoN–DPR(D) 19.01 72.17 70.56 61.09 16.45 56.93 53.99 50.96
-DMoN–DPR(V) 18.89 72.41 69.76 59.10 13.45 59.38 54.03 51.89
-DMoN–DPR(E) 19.73 70.97 71.34 60.40 13.80 60.04 53.17 50.05
-DMoN–DPR(DV) 18.95 72.26 71.33 61.81 13.46 56.85 54.09 56.36
-DMoN–DPR(DE) 19.88 70.56 70.84 61.76 14.22 56.10 54.71 56.94
-DMoN–DPR(VE) 19.70 71.01 71.45 60.56 12.85 55.72 50.14 54.52
-DMoN–DPR(DVE) 19.74 70.68 71.08 62.41 13.46 55.27 54.31 58.86
-5.1 Effectiveness of DMoN–DPR in Clustering
-To assess the effectiveness of DMoN–DPR, we compare the mean values of conductance ($C$), modularity ($Q$), NMI, and pairwise F1 measure across 10 seeds achieved with DMoN–DPR to those obtained with DMoN, as done in (Tsitsulin et al., 2023) (full results can be found in Section B).
-Tables 2 (Cora, CiteSeer, PubMed) and 3 (Coauthor CS, Coauthor Physics) summarize the results, where graph metrics (C, Q) do not rely on ground-truth labels, whereas ‘Labels’ metrics (NMI, F1) compare to known class labels.
-Cora, CiteSeer, and PubMed.
-On Cora, where the diversity of node features and class labels is relatively high, incorporating the distance-based term (D) consistently enhances performance over the baseline DMoN, especially on label-based metrics (NMI and F1).
-The combinations that include the distance term—such as DMoN–DPR(DV), DMoN–DPR(DE), and DMoN–DPR(DVE)—not only improve cluster separability but also boost label-based performance, with DMoN–DPR(DV) and DMoN–DPR(DVE) achieving some of the best F1 and NMI scores.
-The entropy (E) and variance (V) terms alone are less impactful, but when paired with distance, they synergize to produce more balanced and diverse clusters.
-On CiteSeer, which also has a reasonably complex feature space, the introduction of these regularizations yields moderate gains, with combinations that include the entropy term slightly improving label-based metrics.
-In contrast, PubMed, which has a less diverse feature space, sees only marginal benefits.
-Here, the baseline DMoN already performs well, and the added terms provide limited additional advantage.
-Coauthor CS and Coauthor Physics.
-We observe a similar pattern on the two Coauthor datasets.
-Both of these datasets have richer feature spaces than Cora, CiteSeer, and PubMed; however, Coauthor Physics stands out with the largest number of features among all the datasets under consideration (8,415).
-As shown in Table 3, this higher dimensionality helps amplify the benefits of our diversity-preserving regularizations, especially those involving the distance-based term (D).
-Notably, on Coauthor Physics, the pairwise F1 score jumps from 46.80 (DMoN) to 58.86 (DMoN–DPR(DVE)), indicating that combining distance, variance, and entropy can substantially enhance the separability of clusters and the alignment with true labels.
-On Coauthor CS, our proposed regularizations also offer improvements over baseline DMoN, most evident when distance (D) is paired with either variance (V) or entropy (E).
-For example, F1 increases from 58.83 (DMoN) to 62.41 (DMoN–DPR(DVE)).
-Although these gains are meaningful, they are somewhat more moderate than those on Coauthor Physics—likely due to differing distributions of class labels and feature dimensions.
-Overall, these findings suggest that introducing distance-, variance-, and entropy-based regularizations becomes increasingly beneficial as the dimensionality and diversity of node features grow, leading to more pronounced performance boosts on datasets such as Coauthor Physics.
-Analyzing the Influence of Regularization on Clustering Performance.
-Our experiments show that the impact of the distance-based, variance-based, and entropy-based regularizations depends heavily on both the feature richness and label diversity of each dataset.
-On Coauthor CS, and Coauthor Physics, which have relatively higher-dimensional or more varied feature spaces, we observe notable gains in NMI and F1 upon adding these diversity-promoting terms—especially the distance-based (D) regularization, which strongly boosts inter-cluster separation.
-For instance, DMoN–DPR(D) and its variants (e.g., DMoN–DPR(DV), DMoN–DPR(DE), DMoN–DPR(DVE)) yield significant improvements in label alignment on Coauthor Physics, where the F1 score rises by over 10 percentage points compared to baseline DMoN.
-Meanwhile, the variance-based (V) term is helpful in ensuring clusters capture a wider range of node features, an effect that is especially apparent on datasets like Coauthor Physics.
-Although the single V-only variant often yields moderate gains on datasets such as Cora, it generally performs best when combined with the distance or entropy regularizers.
-The entropy-based (E) term aims to produce more balanced cluster distributions, proving modestly beneficial on datasets like CiteSeer and effective on Coauthor CS and Coauthor Physics; however, on PubMed, which has lower feature dimensionality, we see little to no improvement (the numbers remain nearly unchanged from DMoN).
-This indicates that entropy-based balancing may not substantially help in more homogeneous datasets.
-Moreover, combining all three terms (DMoN–DPR(DVE)) often yields the highest F1 scores across multiple datasets (e.g., a leap from 58.83 to 62.41 on Coauthor CS and 46.80 to 58.86 on Coauthor Physics).
-These results suggest that multiple regularizers can work in synergy, but the exact benefits depend on a dataset’s inherent complexity.
-Overall, the findings highlight the importance of tailoring regularization strategies to both the graph structure and feature diversity of the data to achieve the most effective clustering outcomes.
-Trade-Offs Between Structural and Label-Based Metrics
-Our results highlight a trade-off between metrics that prioritize graph structure (e.g., Modularity ($Q$) and Conductance ($C$)) and those that evaluate label alignment (e.g., NMI and F1).
-This trade-off emerges because diversity-preserving regularization terms influence both the feature and label spaces of clusters, leading to distinct outcomes depending on the dataset characteristics and application needs.
-For datasets with rich feature spaces (e.g., Coauthor CS and Coauthor Physics), the diversity-enhancing terms significantly boost label-based metrics like NMI and F1.
-These gains stem from the formation of more homogeneous clusters that align closely with ground-truth labels.
-However, this improvement in label alignment can come at the expense of structural metrics such as modularity, which reward densely connected clusters in the graph topology.
-For instance, while DMoN–DPR(DVE) achieves the highest F1 scores on Coauthor Physics, it exhibits a decrease in modularity compared to baseline DMoN.
-Conversely, on datasets with less diverse or lower-dimensional features (e.g., PubMed), the structural metrics are less affected by the introduction of diversity-preserving regularization terms, and the benefits to label-based metrics are more modest.
-In such cases, the original DMoN, which focuses on structural cohesion, remains a competitive choice for purely unsupervised clustering tasks.
-Ultimately, the decision to prioritize label alignment or structural metrics should be informed by the application context.
-For tasks involving semi-supervised learning or node classification, where accurate label recovery is essential, DMoN–DPR offers clear advantages.
-In contrast, for community detection or unsupervised settings where graph structure is paramount, the baseline DMoN may be preferable.
-This balance between competing objectives underscores the need to tailor clustering approaches to the specific demands of the dataset and task at hand.
-Figure 2: Visualization of clusters (circles) vs.
-ground truth labels (crosses) formed by DMoN on (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics, and by DMoN–DPR on (f) Cora, (g) CiteSeer, (h) PubMed, (i) Coauthor CS, and (j) Coauthor Physics datasets.
-5.2 Visualization of Clusters
-The clustering results visualized in Figure 1 demonstrate that DMoN–DPR generally outperforms DMoN across the benchmark datasets Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics.
-For Cora, DMoN–DPR produces tighter, more well-separated clusters with improved alignment to the ground truth labels, contrasting with the more scattered and less defined clusters formed by DMoN.
-Similarly, for CiteSeer, DMoN–DPR exhibits better cluster compactness and reduced overlap, though the gains are somewhat less pronounced than in Cora.
-On the other hand, PubMed remains challenging for both methods, resulting in notable overlap and scattered clusters; yet, DMoN–DPR still provides slightly tighter groupings.
-Across these three datasets, the diversity-preserving regularization in DMoN–DPR effectively enhances both inter-cluster distinctiveness and intra-cluster variability, producing more meaningful and interpretable partitions.
-Turning to the Coauthor CS and Coauthor Physics datasets, a similar trend emerges.
-While DMoN adequately captures overarching modular structures, it can occasionally struggle to separate nuanced subgroups.
-In contrast, DMoN–DPR consistently yields more cohesive and better-separated clusters, offering clearer decision boundaries and higher alignment with the true labels.
-This improvement is particularly notable in Coauthor CS, where DMoN–DPR visibly reduces overlap and sharpens cluster delineations.
-Although the inherently complex Coauthor Physics dataset can induce some overlap in both methods, DMoN–DPR still produces more robust clusters overall.
-These findings underscore the effectiveness of promoting diversity within hidden representations, as DMoN–DPR not only excels in forming compact and accurate clusters for datasets with greater feature variability but also remains competitive on more challenging datasets.
-5.3 Execution Times and Runtime Analysis
-Table 4 reports the execution times (in seconds) of DMoN and DMoN–DPR on five commonly used graph datasets.
-All experiments were conducted on a CPU to ensure consistent runtime measurement without GPU scheduling effects.
-Table 4: Execution times (in seconds) on CPU for DMoN and DMoN–DPR.
-Dataset DMoN DMoN–DPR
-Cora 32 35
-CiteSeer 76 71
-PubMed 310 300
-Coauthor CS 722 751
-Coauthor Physics 2107 1946
-Although DMoN–DPR includes additional regularization terms for Distance-based, Entropy-based, and Variance-based constraints, its runtime is sometimes comparable to or even marginally lower than plain DMoN.
-The reason is that the most time-consuming part of both methods is dominated by adjacency-based operations (e.g., matrix multiplications with the graph Laplacian or adjacency matrix), which scale on the order of $\mathcal{O}(|E|\,k)$ or $\mathcal{O}(N^{2}k)$ for large, dense graphs (where $N$ is the number of nodes, $E$ is the set of edges, and $k$ is the number of clusters).
-In contrast, the extra DPR calculations—computing entropy across nodes, the variance of cluster assignments, or centroid distances—represent only $\mathcal{O}(N\,k+k^{2}F)$ overhead (with $F$ being the feature dimension) and are typically negligible next to the larger matrix multiplications.
-Furthermore, running on CPU can introduce variability from scheduling, caching, and background processes, so small differences in the recorded times may not be statistically meaningful.
-Thus, although DMoN–DPR performs slightly more computations, the extra cost is overshadowed by the main bottleneck operations, resulting in comparable or even lower measured runtimes in practice.
-5.4 Limitations and Future Directions
-Despite the improvements introduced by our diversity–preserving regularization, DMoN–DPR has a few notable limitations.
-First, tuning the regularization weights for distance, variance, and entropy remains a manual process.
-Determining the best values for these hyperparameters can be time-consuming and sensitive to each dataset’s characteristics.
-An interesting direction for future work is to develop an automated or adaptive scheme that learns these weights during training, making the method more robust and less reliant on exhaustive search.
-Second, while DMoN–DPR yields substantial gains on datasets with a large or diverse feature space, the improvements are often marginal for datasets with lower feature dimensionality (e.g., PubMed).
-Nonetheless, our visualization in Figure 2 suggests that even in these cases, DMoN–DPR tends to produce slightly more coherent cluster structures compared to the baseline, reflecting the benefit of promoting diversity even when numerical metrics (e.g., NMI, F1) show limited change.
-Future research could investigate whether alternative or more specialized regularizers can drive greater numerical improvements on lower-dimensional datasets.
-6 Conclusion
-In this work, we presented DMoN–DPR, an enhanced version of Deep Modularity Networks (DMoN) that incorporates diversity-preserving regularizations to enrich the clustering objective.
-By introducing distance-, entropy-, and variance-based penalties alongside the original modularity and collapse regularization terms, our approach promotes both inter- and intra-cluster diversity.
-Our empirical results on benchmark datasets indicate that DMoN–DPR significantly improves alignment with ground-truth labels, particularly in feature-rich datasets.
-While structural metrics such as modularity and conductance remained competitive, the additional diversity constraints contributed to more interpretable and semantically meaningful cluster formations.
-Overall, DMoN–DPR achieves a balance between interpretability and structural integrity, making it a powerful solution for graph clustering tasks that demand both topological cohesion and semantic differentiation, marking a significant advancement in feature-aware graph clustering.
-Appendix A Ablation Study
-A.1 Varying the Epsilon
+The ablation study is presented in the following subsections.
+D.1.1 Varying the Epsilon
 Figure 2 depicts the ablation study done to find the best value of $\epsilon$ associated with $W_{\text{dist}}$ by setting $W_{\text{var}}$ and $W_{\text{entropy}}$ weights to 0.
@@
 Figure 2: Finding the best value of epsilon ($\epsilon$) by setting $W_{\text{dist}}$ to 1 and $W_{\text{var}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
-A.2 Varying the Variance Weight
+D.1.2 Varying the Variance Weight
 Figure 3 depicts the ablation study done to find the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ weights to 0.
-For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be $1$/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
+For the Cora, CiteSeer, PubMed, Coauthor CS, and Coauthor Physics datasets, the best $W_{\text{var}}$ values were found to be 11/$0.1$, $0.1$, $0.001$, $0.1$, and $1.0$ respectively.
 Figure 3: Finding the best value of $W_{\text{var}}$ by setting $W_{\text{dist}}$ and $W_{\text{entropy}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
-A.3 Varying the Entropy Weight
+D.1.3 Varying the Entropy Weight
 Figure 4 depicts the ablation study done to find the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ weights to 0.
@@
 Figure 4: Finding the best value of $W_{\text{entropy}}$ by setting $W_{\text{dist}}$ and $W_{\text{var}}$ to zero on the (a) Cora, (b) CiteSeer, (c) PubMed, (d) Coauthor CS, and (e) Coauthor Physics datasets.
-Appendix B Full Results
-The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets. The numbers are all in percentage (%).
-B.1 Cora
-The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
+D.1.4 Practical Challenges Regarding Manual Tuning
+We acknowledge that adding diversity‐preserving regularizers introduces more hyperparameters, which can complicate practical adoption. However, in our experiments:
+  • We found that larger coefficients for distance (R_D = 1 or 10) were typically beneficial on highly feature‐rich datasets (Coauthor CS, Coauthor Physics).
+  • Coefficients of 1 to 0.1 for variance (R_V) often worked well across most datasets.
+  • For the entropy term, 0.1 was a good upper bound on large feature spaces, whereas a smaller coefficient was better on less feature‐rich datasets.
+From these observations, we hypothesize that stronger diversity coefficients are advantageous for richer feature spaces, whereas smaller coefficients are sufficient (and sometimes necessary) for lower‐dimensional or simpler datasets.
+As a direction for future work, we suggest exploring automated or adaptive tuning schemes.
+Such an approach could dynamically adjust the weight of each regularizer during training, further reducing manual overhead and improving reproducibility.
+Appendix E Full Results
+The tables below are indicative of the value of each evaluation metric obtained using DMoN/DMoN–DPR pooling for each individual seed for different datasets.
+The seeds were selected at random using a random number generator.
+The numbers are all in percentage (%).
+E.1 Cora
+The following tables list the performance of DiffPool, MinCut Pool, DMoN and different DMoN–DPR configurations across different seeds on the Cora dataset.
+Table 8: Results obtained by using DiffPool on the Cora dataset.
+Seed Method Conductance Modularity NMI F1
+993 DiffPool 15.90 63.81 42.06 51.72
+550 DiffPool 18.74 61.16 35.22 36.23
+243 DiffPool 15.84 64.02 40.57 49.64
+16 DiffPool 17.07 57.90 37.43 46.73
+716 DiffPool 15.33 64.34 46.69 53.05
+383 DiffPool 15.54 63.93 39.65 44.05
+277 DiffPool 12.83 67.00 47.31 55.25
+274 DiffPool 16.31 63.31 36.01 39.97
+188 DiffPool 18.26 56.66 31.67 35.33
+796 DiffPool 14.08 65.66 44.65 53.50
+Mean $\pm$ Std 15.99 $\pm$ 1.77 62.78 $\pm$ 3.28 40.13 $\pm$ 5.15 46.55 $\pm$ 7.34
+Table 9: Results obtained by using MinCut pooling on the Cora dataset.
+Seed Method Conductance Modularity NMI F1
+993 MinCut 12.73 72.59 41.66 43.90
+550 MinCut 11.61 73.70 45.63 45.03
+243 MinCut 13.89 71.56 31.48 33.71
+16 MinCut 14.08 71.37 38.06 37.21
+716 MinCut 14.38 71.14 33.82 40.37
+383 MinCut 15.18 70.36 29.04 28.48
+277 MinCut 13.09 72.42 42.81 46.15
+274 MinCut 13.32 72.16 39.60 39.77
+188 MinCut 13.81 71.65 40.52 44.13
+796 MinCut 14.44 70.91 34.75 32.71
+Mean $\pm$ Std 13.65 $\pm$ 1.01 71.79 $\pm$ 0.96 37.74 $\pm$ 5.31 39.15 $\pm$ 5.98
 Table 10: Results obtained by using DMoN pooling on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN 10.25 73.93 48.71 55.38
-550 DMoN 9.95 71.77 40.62 42.17
-243 DMoN 10.06 73.74 47.62 53.20
-16 DMoN 12.13 70.62 44.77 46.61
-716 DMoN 11.92 70.40 37.87 39.65
-383 DMoN 11.20 71.80 44.40 52.46
-277 DMoN 11.03 72.50 37.07 36.07
-274 DMoN 10.12 74.14 45.30 44.74
-188 DMoN 11.48 73.36 46.79 49.47
-796 DMoN 11.86 71.67 43.30 42.69
-Mean 11.00 72.39 43.65 46.24
-Variance 0.01 0.02 0.16 0.40
+993 DMoN–DPR (DVE) 11.35 71.32 48.90 52.56
+550 DMoN–DPR (DVE) 9.81 72.13 38.15 42.84
+243 DMoN–DPR (DVE) 10.72 73.08 40.71 40.03
+16 DMoN–DPR (DVE) 11.06 71.48 42.84 46.41
+716 DMoN–DPR (DVE) 9.93 72.55 42.77 44.43
+383 DMoN–DPR (DVE) 11.54 73.41 44.12 51.88
+277 DMoN–DPR (DVE) 9.66 73.52 45.61 49.43
+274 DMoN–DPR (DVE) 11.03 74.44 41.98 41.83
+188 DMoN–DPR (DVE) 9.51 74.37 51.31 56.52
+796 DMoN–DPR (DVE) 10.65 71.36 42.78 43.37
+Mean $\pm$ Std 10.53 $\pm$ 0.74 72.77 $\pm$ 1.18 43.92 $\pm$ 3.85 46.93 $\pm$ 5.42
 Table 11: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, and distance weight of 1 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (D) 10.50 73.78 51.01 57.45
-550 DMoN–DPR (D) 10.06 72.20 41.49 44.43
-243 DMoN–DPR (D) 9.49 73.58 47.80 54.94
-16 DMoN–DPR (D) 10.55 72.20 46.22 49.19
-716 DMoN–DPR (D) 10.38 71.10 39.29 43.43
-383 DMoN–DPR (D) 9.49 74.11 48.02 54.77
-277 DMoN–DPR (D) 10.65 72.45 40.12 42.77
-274 DMoN–DPR (D) 9.47 74.41 47.64 52.47
-188 DMoN–DPR (D) 10.44 73.75 48.41 54.78
-796 DMoN–DPR (D) 11.29 72.13 43.05 45.75
-Mean 10.23 72.97 45.31 50.00
-Variance 0.00 0.01 0.16 0.31
+993 DMoN–DPR (D) 11.31 71.37 49.62 53.10
+550 DMoN–DPR (D) 9.76 72.28 40.96 45.72
+243 DMoN–DPR (D) 10.40 74.23 43.14 44.89
+16 DMoN–DPR (D) 11.90 71.00 42.22 44.14
+716 DMoN–DPR (D) 9.62 71.51 43.06 46.72
+383 DMoN–DPR (D) 11.24 72.86 43.27 51.09
+277 DMoN–DPR (D) 13.64 69.81 40.51 44.53
+274 DMoN–DPR (D) 11.12 73.24 43.58 45.78
+188 DMoN–DPR (D) 9.80 74.15 50.89 56.19
+796 DMoN–DPR (D) 10.70 71.52 42.57 42.95
+Mean $\pm$ Std 10.95 $\pm$ 1.21 72.20 $\pm$ 1.42 43.98 $\pm$ 3.46 47.51 $\pm$ 4.40
 Table 12: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR 10.99 72.41 51.86 58.92
-550 DMoN–DPR (V) 11.10 70.31 40.04 43.09
-243 DMoN–DPR (V) 11.75 71.02 47.47 54.54
-16 DMoN–DPR (V) 10.99 70.18 45.82 48.81
-716 DMoN–DPR (V) 11.08 69.84 40.82 41.88
-383 DMoN–DPR (V) 12.13 69.37 42.89 51.12
-277 DMoN–DPR (V) 11.42 70.91 37.43 37.64
-274 DMoN–DPR (V) 11.05 72.61 46.97 48.40
-188 DMoN–DPR (V) 13.11 70.74 45.71 52.26
-796 DMoN–DPR (V) 12.79 70.56 43.70 45.13
-Mean 11.64 70.80 44.27 48.18
-Variance 0.01 0.01 0.18 0.41
+993 DMoN–DPR (V) 11.25 70.12 48.19 52.60
+550 DMoN–DPR (V) 10.95 70.66 39.21 43.64
+243 DMoN–DPR (V) 11.63 71.41 40.33 39.36
+16 DMoN–DPR (V) 11.25 70.02 43.27 45.92
+716 DMoN–DPR (V) 11.22 70.58 43.18 45.43
+383 DMoN–DPR (V) 13.40 70.06 42.14 51.60
+277 DMoN–DPR (V) 11.33 69.87 41.46 45.44
+274 DMoN–DPR (V) 12.43 72.57 45.58 44.85
+188 DMoN–DPR (V) 12.83 70.46 46.59 51.89
+796 DMoN–DPR (V) 12.37 69.61 43.49 43.14
+Mean $\pm$ Std 11.87 $\pm$ 0.83 70.54 $\pm$ 0.88 43.34 $\pm$ 2.79 46.39 $\pm$ 4.32
 Table 13: Results obtained by using DMoN–DPR pooling, using the best entropy weight of 0.001 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (E) 10.25 73.81 49.10 55.86
-550 DMoN–DPR (E) 10.06 71.70 40.49 42.06
-243 DMoN–DPR (E) 9.91 73.82 48.60 53.72
-16 DMoN–DPR (E) 12.32 70.46 44.95 46.88
-716 DMoN–DPR (E) 11.86 70.43 37.72 39.66
-383 DMoN–DPR (E) 11.01 71.93 44.49 52.83
-277 DMoN–DPR (E) 11.16 72.37 37.38 35.98
-274 DMoN–DPR (E) 10.17 74.20 44.13 44.02
-188 DMoN–DPR (E) 11.41 73.37 46.28 48.86
-796 DMoN–DPR (E) 11.88 71.65 43.17 42.63
-Mean 11.00 72.37 43.63 46.25
-Variance 0.01 0.02 0.16 0.43
+993 DMoN–DPR (E) 11.20 71.38 48.47 52.36
+550 DMoN–DPR (E) 9.91 72.05 38.08 42.93
+243 DMoN–DPR (E) 10.57 73.23 40.85 39.97
+16 DMoN–DPR (E) 11.06 71.47 42.93 46.56
+716 DMoN–DPR (E) 9.89 72.54 42.62 44.72
+383 DMoN–DPR (E) 11.42 73.51 44.30 51.74
+277 DMoN–DPR (E) 9.55 73.52 45.49 49.53
+274 DMoN–DPR (E) 11.08 74.33 43.10 42.42
+188 DMoN–DPR (E) 9.61 74.26 51.24 56.67
+796 DMoN–DPR (E) 10.63 71.45 42.93 43.32
+Mean $\pm$ Std 10.49 $\pm$ 0.70 72.77 $\pm$ 1.15 44.00 $\pm$ 3.72 47.02 $\pm$ 5.34
 Table 14: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DV) 10.25 73.42 51.71 58.37
-550 DMoN–DPR (DV) 10.17 72.08 41.28 44.08
-243 DMoN–DPR (DV) 9.66 73.18 48.08 55.94
-16 DMoN–DPR (DV) 10.72 71.69 46.80 49.51
-716 DMoN–DPR (DV) 10.52 70.97 39.46 43.32
-383 DMoN–DPR (DV) 9.87 73.82 48.11 54.87
-277 DMoN–DPR (DV) 10.95 71.90 39.01 42.43
-274 DMoN–DPR (DV) 9.61 74.13 48.04 53.26
-188 DMoN–DPR (DV) 10.57 73.51 48.47 54.55
-796 DMoN–DPR (DV) 11.39 72.06 43.12 44.93
-Mean 10.37 72.68 45.41 50.13
-Variance 0.00 0.01 0.19 0.36
+993 DMoN–DPR (DV) 10.93 71.05 49.87 53.66
+550 DMoN–DPR (DV) 9.51 72.33 41.04 45.65
+243 DMoN–DPR (DV) 10.84 73.28 42.18 42.58
+16 DMoN–DPR (DV) 10.86 71.23 43.08 44.77
+716 DMoN–DPR (DV) 9.61 71.39 44.18 47.59
+383 DMoN–DPR (DV) 11.80 73.16 43.66 46.75
+277 DMoN–DPR (DV) 13.74 69.61 40.23 44.66
+274 DMoN–DPR (DV) 11.33 73.29 45.41 47.47
+188 DMoN–DPR (DV) 9.61 74.03 51.32 56.96
+796 DMoN–DPR (DV) 10.72 71.30 43.02 43.54
+Mean $\pm$ Std 10.89 $\pm$ 1.26 72.07 $\pm$ 1.37 44.40 $\pm$ 3.60 47.36 $\pm$ 4.55
 Table 15: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.001 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DE) 10.63 73.82 51.30 57.10
-550 DMoN–DPR (DE) 10.10 72.15 41.06 44.04
-243 DMoN–DPR (DE) 9.68 73.25 47.62 55.30
-16 DMoN–DPR (DE) 10.61 72.13 46.23 49.09
-716 DMoN–DPR (DE) 10.63 70.90 39.37 43.28
-383 DMoN–DPR (DE) 9.53 74.04 47.93 54.43
-277 DMoN–DPR (DE) 10.89 72.25 39.14 42.53
-274 DMoN–DPR (DE) 9.49 74.39 48.31 52.84
-188 DMoN–DPR (DE) 10.52 73.68 48.16 54.61
-796 DMoN–DPR (DE) 11.18 72.21 42.71 45.56
-Mean 10.33 72.88 45.18 49.88
-Variance 0.00 0.01 0.18 0.32
+993 DMoN–DPR (DE) 11.16 71.38 49.75 53.29
+550 DMoN–DPR (DE) 9.68 72.33 40.85 45.67
+243 DMoN–DPR (DE) 10.08 74.45 43.94 46.11
+16 DMoN–DPR (DE) 11.29 71.57 43.55 43.97
+716 DMoN–DPR (DE) 9.61 71.58 44.19 47.71
+383 DMoN–DPR (DE) 11.73 73.06 42.66 48.52
+277 DMoN–DPR (DE) 13.58 69.91 40.43 44.48
+274 DMoN–DPR (DE) 11.01 73.29 43.61 45.77
+188 DMoN–DPR (DE) 9.81 74.04 51.34 56.67
+796 DMoN–DPR (DE) 10.82 71.36 42.67 43.10
+Mean $\pm$ Std 10.88 $\pm$ 1.21 72.30 $\pm$ 1.40 44.30 $\pm$ 3.54 47.53 $\pm$ 4.32
 Table 16: Results obtained by using DMoN–DPR pooling, using the best variance weight of 1, and entropy weight of 0.001 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (VE) 10.86 72.34 51.43 58.35
-550 DMoN–DPR (VE) 11.10 70.30 40.10 43.20
-243 DMoN–DPR (VE) 11.71 71.05 47.46 54.42
-16 DMoN–DPR (VE) 11.06 70.19 45.89 48.85
-716 DMoN–DPR (VE) 11.01 69.83 40.77 41.82
-383 DMoN–DPR (VE) 12.14 69.28 42.65 50.92
-277 DMoN–DPR (VE) 11.46 70.98 37.40 37.53
-274 DMoN–DPR (VE) 11.14 72.47 46.72 48.43
-188 DMoN–DPR (VE) 13.11 70.78 45.36 52.19
-796 DMoN–DPR (VE) 12.85 70.48 43.74 45.16
-Mean 11.64 70.77 44.15 48.09
-Variance 0.01 0.01 0.17 0.39
+993 DMoN–DPR (VE) 11.16 70.08 48.03 52.85
+550 DMoN–DPR (VE) 10.86 70.77 39.39 43.76
+243 DMoN–DPR (VE) 11.78 71.27 40.17 39.23
+16 DMoN–DPR (VE) 11.29 69.93 43.33 45.86
+716 DMoN–DPR (VE) 11.10 70.70 43.42 45.66
+383 DMoN–DPR (VE) 13.17 70.14 41.54 51.22
+277 DMoN–DPR (VE) 11.46 69.60 41.28 45.35
+274 DMoN–DPR (VE) 12.43 72.56 45.66 44.94
+188 DMoN–DPR (VE) 12.90 70.39 46.46 51.76
+796 DMoN–DPR (VE) 12.41 69.61 43.10 42.98
+Mean $\pm$ Std 11.86 $\pm$ 0.82 70.50 $\pm$ 0.89 43.24 $\pm$ 2.79 46.36 $\pm$ 4.32
 Table 17: Results obtained by using DMoN–DPR pooling, using the best epsilon value of 0.0001, variance weight of 0.1, and entropy weight of 0.001 on the Cora dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DVE) 10.06 73.55 51.62 58.48
-550 DMoN–DPR (DVE) 10.16 72.04 41.47 44.30
-243 DMoN–DPR (DVE) 9.87 72.96 47.86 55.57
-16 DMoN–DPR (DVE) 10.76 71.64 46.63 49.25
-716 DMoN–DPR (DVE) 10.72 70.81 39.64 43.32
-383 DMoN–DPR (DVE) 9.93 73.75 48.14 54.86
-277 DMoN–DPR (DVE) 10.89 71.99 39.10 42.50
-274 DMoN–DPR (DVE) 9.83 73.94 47.81 52.85
-188 DMoN–DPR (DVE) 10.42 73.73 48.53 54.50
-796 DMoN–DPR (DVE) 11.41 72.05 43.51 44.96
-Mean 10.41 72.65 45.43 50.06
-Variance 0.00 0.01 0.18 0.35
-B.2 CiteSeer
-The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
+993 DMoN–DPR (DVE) 10.65 71.20 49.95 53.90
+550 DMoN–DPR (DVE) 9.51 72.33 41.04 45.64
+243 DMoN–DPR (DVE) 10.86 73.34 41.92 42.43
+16 DMoN–DPR (DVE) 10.78 71.29 42.83 44.73
+716 DMoN–DPR (DVE) 9.74 71.30 43.98 47.36
+383 DMoN–DPR (DVE) 12.58 72.50 42.83 45.78
+277 DMoN–DPR (DVE) 13.81 69.51 40.61 45.21
+274 DMoN–DPR (DVE) 10.97 73.71 46.42 47.92
+188 DMoN–DPR (DVE) 9.64 73.99 51.33 56.97
+796 DMoN–DPR (DVE) 10.80 71.35 42.74 43.29
+Mean $\pm$ Std 10.93 $\pm$ 1.34 72.05 $\pm$ 1.39 44.37 $\pm$ 3.69 47.32 $\pm$ 4.63
+E.2 CiteSeer
+The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the CiteSeer dataset.
+Table 18: Results obtained by using DiffPool pooling on the CiteSeer dataset.
+Seed Method Conductance Modularity NMI F1
+993 DiffPool 8.24 64.76 33.99 48.98
+550 DiffPool 7.27 70.84 38.27 51.01
+243 DiffPool 6.59 66.78 34.97 51.19
+16 DiffPool 10.43 68.61 34.15 47.16
+716 DiffPool 7.01 70.87 33.12 46.30
+383 DiffPool 9.49 56.09 28.78 45.91
+277 DiffPool 9.58 64.40 29.61 41.80
+274 DiffPool 6.50 71.68 32.23 44.80
+188 DiffPool 7.18 66.15 32.83 48.86
+796 DiffPool 6.94 66.72 36.02 52.33
+Mean $\pm$ Std 7.92 $\pm$ 1.42 66.69 $\pm$ 4.53 33.40 $\pm$ 2.82 47.83 $\pm$ 3.27
+Table 19: Results obtained by using MinCut pooling on the CiteSeer dataset.
+Seed Method Conductance Modularity NMI F1
+993 MinCut 5.45 76.59 33.17 45.03
+550 MinCut 5.34 75.36 24.92 35.92
+243 MinCut 5.01 76.67 30.26 42.82
+16 MinCut 7.21 73.20 18.09 26.08
+716 MinCut 7.01 74.82 22.11 30.09
+383 MinCut 6.04 75.25 23.28 34.72
+277 MinCut 7.05 73.86 17.40 25.14
+274 MinCut 5.32 76.16 25.47 38.90
+188 MinCut 6.44 75.23 32.94 42.54
+796 MinCut 7.05 74.93 24.43 31.51
+Mean $\pm$ Std 6.19 $\pm$ 0.86 75.21 $\pm$ 1.11 25.21 $\pm$ 5.52 35.28 $\pm$ 7.05
 Table 20: Results obtained by using DMoN pooling on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN 6.09 74.10 22.56 26.60
-550 DMoN 4.31 75.22 25.92 41.43
-243 DMoN 4.90 78.26 34.20 46.62
-16 DMoN 5.87 75.48 28.89 38.28
-716 DMoN 3.60 76.57 38.40 52.01
-383 DMoN 4.94 75.34 26.50 39.64
-277 DMoN 6.50 75.67 28.87 37.96
-274 DMoN 4.35 74.93 32.36 47.53
-188 DMoN 4.50 76.41 34.86 46.98
-796 DMoN 4.13 75.24 30.07 44.67
-Mean 4.92 75.72 30.26 42.17
-Variance 0.01 0.01 0.23 0.51
+993 DMoN 6.81 74.92 20.78 26.09
+550 DMoN 4.11 75.53 26.91 42.64
+243 DMoN 4.28 75.57 32.12 45.74
+16 DMoN 4.64 77.22 33.11 44.56
+716 DMoN 4.53 76.24 36.95 50.12
+383 DMoN 4.24 75.55 33.25 47.00
+277 DMoN 4.26 75.63 29.25 42.33
+274 DMoN 4.48 75.77 30.83 45.65
+188 DMoN 6.17 72.74 24.52 36.01
+796 DMoN 5.12 75.34 28.98 44.42
+Mean $\pm$ Std 4.86 $\pm$ 0.91 75.45 $\pm$ 1.13 29.67 $\pm$ 4.70 42.46 $\pm$ 6.82
 Table 21: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, and distance weight of 1 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (D) 6.30 74.07 22.96 27.06
-550 DMoN–DPR (D) 4.11 75.45 26.52 42.31
-243 DMoN–DPR (D) 3.98 76.17 36.63 51.64
-16 DMoN–DPR (D) 5.36 76.07 29.32 42.93
-716 DMoN–DPR (D) 4.17 76.07 39.05 51.65
-383 DMoN–DPR (D) 5.18 75.60 26.83 37.54
-277 DMoN–DPR (D) 5.16 75.30 29.25 40.17
-274 DMoN–DPR (D) 4.28 75.19 32.02 46.99
-188 DMoN–DPR (D) 4.70 76.60 34.34 45.97
-796 DMoN–DPR (D) 4.24 75.92 29.15 41.29
-Mean 4.75 75.64 30.61 42.76
-Variance 0.01 0.00 0.24 0.52
+993 DMoN–DPR (D) 7.18 74.27 23.90 27.73
+550 DMoN–DPR (D) 4.22 75.80 26.23 41.98
+243 DMoN–DPR (D) 3.82 76.02 35.91 50.13
+16 DMoN–DPR (D) 4.83 76.31 32.78 44.47
+716 DMoN–DPR (D) 4.57 76.13 36.99 50.14
+383 DMoN–DPR (D) 4.81 75.97 33.02 45.96
+277 DMoN–DPR (D) 3.87 75.75 29.37 42.34
+274 DMoN–DPR (D) 4.15 76.06 30.88 45.48
+188 DMoN–DPR (D) 6.37 73.98 25.00 36.47
+796 DMoN–DPR (D) 5.71 75.26 26.83 39.93
+Mean $\pm$ Std 4.95 $\pm$ 1.12 75.55 $\pm$ 0.81 30.09 $\pm$ 4.57 42.46 $\pm$ 6.69
 Table 22: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (V) 6.33 73.98 23.33 26.50
-550 DMoN–DPR (V) 4.17 75.20 26.64 42.18
-243 DMoN–DPR (V) 4.99 78.15 34.08 46.34
-16 DMoN–DPR (V) 6.02 75.21 29.38 38.80
-716 DMoN–DPR (V) 3.71 76.45 38.70 52.38
-383 DMoN–DPR (V) 5.16 74.60 26.80 39.65
-277 DMoN–DPR (V) 5.84 75.26 29.13 39.42
-274 DMoN–DPR (V) 4.42 74.79 31.87 46.92
-188 DMoN–DPR (V) 4.33 76.39 34.76 47.14
-796 DMoN–DPR (V) 4.33 75.87 29.92 43.30
-Mean 4.93 75.59 30.46 42.26
-Variance 0.01 0.01 0.20 0.49
+993 DMoN–DPR (V) 6.83 74.31 21.24 28.03
+550 DMoN–DPR (V) 4.04 75.60 26.97 42.75
+243 DMoN–DPR (V) 4.28 75.41 32.23 45.77
+16 DMoN–DPR (V) 5.14 76.33 34.01 46.10
+716 DMoN–DPR (V) 4.57 76.13 36.89 50.11
+383 DMoN–DPR (V) 4.15 75.38 33.48 47.59
+277 DMoN–DPR (V) 4.35 75.42 28.95 42.12
+274 DMoN–DPR (V) 4.59 75.13 30.92 46.21
+188 DMoN–DPR (V) 6.39 73.03 24.67 35.88
+796 DMoN–DPR (V) 5.21 74.93 28.27 43.93
+Mean $\pm$ Std 4.96 $\pm$ 0.96 75.17 $\pm$ 0.94 29.76 $\pm$ 4.71 42.85 $\pm$ 6.46
 Table 23: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.01 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (E) 6.59 74.00 22.17 27.07
-550 DMoN–DPR (E) 4.42 75.06 26.32 41.61
-243 DMoN–DPR (E) 5.14 77.95 34.48 46.79
-16 DMoN–DPR (E) 6.15 75.09 29.46 38.54
-716 DMoN–DPR (E) 3.67 76.36 39.12 52.80
-383 DMoN–DPR (E) 5.32 74.52 27.05 39.70
-277 DMoN–DPR (E) 5.89 75.27 29.14 39.34
-274 DMoN–DPR (E) 4.35 74.85 32.18 47.28
-188 DMoN–DPR (E) 4.55 76.16 34.69 47.23
-796 DMoN–DPR (E) 4.11 75.02 30.58 45.06
-Mean 5.02 75.43 30.52 42.54
-Variance 0.01 0.01 0.24 0.50
+993 DMoN–DPR (E) 7.27 73.95 21.36 28.28
+550 DMoN–DPR (E) 4.15 75.67 26.79 42.59
+243 DMoN–DPR (E) 4.31 75.39 32.27 45.75
+16 DMoN–DPR (E) 4.92 76.51 34.13 46.04
+716 DMoN–DPR (E) 4.48 76.00 37.03 50.46
+383 DMoN–DPR (E) 4.46 75.36 32.82 46.38
+277 DMoN–DPR (E) 4.20 75.58 29.78 43.08
+274 DMoN–DPR (E) 4.26 75.81 31.63 46.59
+188 DMoN–DPR (E) 6.22 73.30 25.10 36.09
+796 DMoN–DPR (E) 5.07 74.98 28.59 44.38
+Mean $\pm$ Std 4.93 $\pm$ 1.03 75.25 $\pm$ 0.96 29.95 $\pm$ 4.63 42.96 $\pm$ 6.35
 Table 24: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and variance weight of 0.1 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DV) 6.33 73.81 23.64 27.24
-550 DMoN–DPR (DV) 4.09 75.48 26.65 42.08
-243 DMoN–DPR (DV) 3.89 76.17 35.67 51.00
-16 DMoN–DPR (DV) 5.71 75.73 29.33 42.91
-716 DMoN–DPR (DV) 4.11 76.03 38.91 51.73
-383 DMoN–DPR (DV) 5.16 75.31 27.06 38.22
-277 DMoN–DPR (DV) 5.01 75.32 29.81 40.68
-274 DMoN–DPR (DV) 4.35 74.92 32.29 47.22
-188 DMoN–DPR (DV) 4.50 76.36 34.12 46.47
-796 DMoN–DPR (DV) 4.44 75.93 28.46 40.94
-Mean 4.76 75.51 30.59 42.85
-Variance 0.01 0.01 0.21 0.50
+993 DMoN–DPR 7.54 73.39 23.91 29.07
+550 DMoN–DPR 4.35 75.62 26.61 42.29
+243 DMoN–DPR 3.84 75.78 36.17 50.51
+16 DMoN–DPR 4.86 76.09 33.03 44.97
+716 DMoN–DPR 4.55 76.03 38.13 51.55
+383 DMoN–DPR 4.90 75.53 33.12 46.09
+277 DMoN–DPR 4.06 75.33 29.89 42.79
+274 DMoN–DPR 4.28 75.82 31.51 46.34
+188 DMoN–DPR 6.50 73.71 25.38 36.63
+796 DMoN–DPR 6.04 74.53 27.25 42.22
+Mean $\pm$ Std 5.09 $\pm$ 1.21 75.18 $\pm$ 0.97 30.50 $\pm$ 4.72 43.25 $\pm$ 6.58
 Table 25: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, and entropy weight of 0.01 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DE) 6.48 73.71 23.56 27.55
-550 DMoN–DPR (DE) 4.31 75.24 26.29 41.96
-243 DMoN–DPR (DE) 3.95 76.11 36.10 51.46
-16 DMoN–DPR (DE) 5.76 75.68 29.78 43.37
-716 DMoN–DPR (DE) 4.15 76.00 39.03 51.88
-383 DMoN–DPR (DE) 5.43 75.09 27.51 38.72
-277 DMoN–DPR (DE) 5.10 75.25 29.91 40.77
-274 DMoN–DPR (DE) 4.42 74.89 31.85 46.74
-188 DMoN–DPR (DE) 4.75 76.28 34.05 46.33
-796 DMoN–DPR (DE) 4.26 75.21 29.63 43.94
-Mean 4.86 75.35 30.77 43.27
-Variance 0.01 0.01 0.22 0.49
+993 DMoN–DPR (DE) 7.45 73.34 24.03 29.50
+550 DMoN–DPR (DE) 4.37 75.63 26.92 42.46
+243 DMoN–DPR (DE) 3.89 75.81 36.33 50.53
+16 DMoN–DPR (DE) 5.07 75.80 32.09 43.42
+716 DMoN–DPR (DE) 4.33 75.97 38.51 52.05
+383 DMoN–DPR (DE) 4.83 75.58 32.59 45.72
+277 DMoN–DPR (DE) 4.02 75.47 30.22 43.10
+274 DMoN–DPR (DE) 4.35 75.80 31.50 46.10
+188 DMoN–DPR (DE) 6.35 73.93 25.29 36.35
+796 DMoN–DPR (DE) 6.06 74.49 27.23 42.08
+Mean $\pm$ Std 5.07 $\pm$ 1.17 75.18 $\pm$ 0.92 30.47 $\pm$ 4.70 43.13 $\pm$ 6.53
 Table 26: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and best entropy weight of 0.01 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (VE) 6.50 73.55 22.91 27.80
-550 DMoN–DPR (VE) 4.59 74.88 26.41 41.35
-243 DMoN–DPR (VE) 5.25 77.84 34.60 47.05
-16 DMoN–DPR (VE) 5.95 76.05 31.65 40.71
-716 DMoN–DPR (VE) 3.84 76.18 39.28 52.93
-383 DMoN–DPR (VE) 5.40 74.12 26.44 39.40
-277 DMoN–DPR (VE) 5.91 75.11 29.14 39.27
-274 DMoN–DPR (VE) 4.55 74.52 32.02 47.12
-188 DMoN–DPR (VE) 4.70 75.85 34.92 47.37
-796 DMoN–DPR (VE) 4.24 74.50 31.26 45.61
-Mean 5.09 75.26 30.86 42.86
-Variance 0.01 0.02 0.23 0.47
+993 DMoN–DPR (VE) 7.51 73.51 21.85 28.66
+550 DMoN–DPR (VE) 4.35 75.43 26.50 42.19
+243 DMoN–DPR (VE) 4.48 75.15 32.58 46.11
+16 DMoN–DPR (VE) 4.79 76.51 33.09 45.03
+716 DMoN–DPR (VE) 4.50 75.95 37.46 51.15
+383 DMoN–DPR (VE) 4.61 74.94 33.20 46.98
+277 DMoN–DPR (VE) 4.24 75.33 30.08 43.40
+274 DMoN–DPR (VE) 4.37 75.69 31.72 46.90
+188 DMoN–DPR (VE) 6.39 72.92 25.40 36.39
+796 DMoN–DPR (VE) 5.23 74.71 29.84 45.84
+Mean $\pm$ Std 5.05 $\pm$ 1.07 75.01 $\pm$ 1.09 30.17 $\pm$ 4.53 43.27 $\pm$ 6.40
 Table 27: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.0001, distance weight of 1, variance weight of 0.1, and entropy weight of 0.01 on the CiteSeer dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DVE) 6.39 73.36 23.60 28.45
-550 DMoN–DPR (DVE) 4.35 75.08 26.35 41.77
-243 DMoN–DPR (DVE) 3.93 75.80 36.58 52.14
-16 DMoN–DPR (DVE) 5.38 76.03 28.50 41.56
-716 DMoN–DPR (DVE) 4.22 75.83 38.91 51.92
-383 DMoN–DPR (DVE) 5.34 74.92 28.02 39.07
-277 DMoN–DPR (DVE) 5.34 75.06 29.87 40.58
-274 DMoN–DPR (DVE) 4.46 74.65 31.76 46.98
-188 DMoN–DPR (DVE) 4.68 75.90 34.46 47.48
-796 DMoN–DPR (DVE) 4.44 74.79 29.69 44.27
-Mean 4.85 75.14 30.77 43.42
-Variance 0.01 0.01 0.22 0.48
-B.3 PubMed
-The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.
+993 DMoN–DPR 7.45 73.17 23.84 29.48
+550 DMoN–DPR 4.15 75.58 26.66 42.48
+243 DMoN–DPR 3.91 75.68 36.30 50.67
+16 DMoN–DPR 5.23 75.44 31.58 42.92
+716 DMoN–DPR 4.59 75.96 37.14 50.08
+383 DMoN–DPR 4.75 75.43 32.95 45.93
+277 DMoN–DPR 4.44 75.10 29.34 41.94
+274 DMoN–DPR 4.35 75.91 31.30 46.10
+188 DMoN–DPR 6.52 73.57 26.09 37.49
+796 DMoN–DPR 6.24 74.21 27.32 42.91
+Mean $\pm$ Std 5.16 $\pm$ 1.18 75.01 $\pm$ 1.00 30.25 $\pm$ 4.41 43.00 $\pm$ 6.16
+E.3 PubMed
+The following tables list the performance of MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the PubMed dataset.
+Table 28: Results obtained by using MinCut pooling on the PubMed dataset.
+Seed Method Conductance Modularity NMI F1
+993 MinCut 12.82 53.38 23.17 41.29
+550 MinCut 8.41 56.48 21.15 44.79
+243 MinCut 12.52 53.77 23.64 40.72
+16 MinCut 12.60 53.65 23.65 39.98
+716 MinCut 10.43 55.36 21.44 37.44
+383 MinCut 12.62 53.67 23.85 41.18
+277 MinCut 8.32 56.56 20.47 44.08
+274 MinCut 8.33 56.46 20.25 44.46
+188 MinCut 12.52 53.76 23.34 38.89
+796 MinCut 12.79 53.53 23.84 40.30
+Mean $\pm$ Std 11.14 $\pm$ 2.04 54.66 $\pm$ 1.38 22.48 $\pm$ 1.47 41.31 $\pm$ 2.44
 Table 29: Results obtained by using DMoN pooling on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN 10.64 55.82 21.92 39.57
-550 DMoN 7.91 58.30 26.08 48.88
-243 DMoN 7.87 57.66 21.90 44.67
-16 DMoN 8.02 57.40 21.90 44.83
-716 DMoN 8.02 57.35 21.98 45.16
-383 DMoN 12.56 53.65 26.03 49.14
-277 DMoN 7.78 57.61 21.73 44.90
-274 DMoN 7.94 57.44 21.59 44.78
-188 DMoN 7.90 57.50 21.86 45.03
-796 DMoN 7.93 57.53 21.76 44.82
-Mean 8.66 57.03 22.68 45.18
-Variance 0.03 0.02 0.03 0.07
+993 DMoN 10.94 55.54 22.49 40.20
+550 DMoN 7.64 58.82 26.92 49.53
+243 DMoN 7.93 57.62 21.77 44.56
+16 DMoN 7.96 57.63 21.74 44.37
+716 DMoN 8.05 57.34 21.70 44.97
+383 DMoN 9.79 56.12 22.79 37.01
+277 DMoN 7.80 57.63 21.69 44.89
+274 DMoN 7.90 57.45 21.60 44.80
+188 DMoN 10.09 55.60 21.39 37.09
+796 DMoN 7.96 57.52 21.76 44.85
+Mean $\pm$ Std 8.61 $\pm$ 1.19 57.13 $\pm$ 1.04 22.38 $\pm$ 1.65 43.23 $\pm$ 3.93
 Table 30: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001 and distance weight of 1 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (D) 10.63 55.83 21.92 39.59
-550 DMoN–DPR (D) 7.94 58.27 26.16 48.94
-243 DMoN–DPR (D) 7.88 57.65 21.93 44.67
-16 DMoN–DPR (D) 8.00 57.41 21.80 44.75
-716 DMoN–DPR (D) 8.07 57.30 21.97 45.11
-383 DMoN–DPR (D) 12.53 53.69 25.84 49.03
-277 DMoN–DPR (D) 7.79 57.60 21.73 44.91
-274 DMoN–DPR (D) 7.96 57.43 21.76 44.94
-188 DMoN–DPR (D) 7.91 57.49 21.87 45.05
-796 DMoN–DPR (D) 7.92 57.54 21.84 44.89
-Mean 8.66 57.02 22.68 45.19
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (D) 10.82 55.67 22.35 40.01
+550 DMoN–DPR (D) 7.65 58.81 26.89 49.57
+243 DMoN–DPR (D) 7.95 57.60 21.79 44.56
+16 DMoN–DPR (D) 7.99 57.59 21.73 44.35
+716 DMoN–DPR (D) 8.05 57.36 21.71 44.95
+383 DMoN–DPR (D) 9.78 56.13 22.83 37.08
+277 DMoN–DPR (D) 7.81 57.62 21.67 44.88
+274 DMoN–DPR (D) 7.82 57.53 21.67 44.91
+188 DMoN–DPR (D) 10.14 55.54 21.34 37.09
+796 DMoN–DPR (D) 7.97 57.51 21.82 44.96
+Mean $\pm$ Std 8.60 $\pm$ 1.17 57.14 $\pm$ 1.03 22.38 $\pm$ 1.64 43.24 $\pm$ 3.95
 Table 31: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (V) 10.64 55.82 21.91 39.57
-550 DMoN–DPR (V) 7.92 58.29 26.08 48.88
-243 DMoN–DPR (V) 7.88 57.65 21.91 44.68
-16 DMoN–DPR (V) 8.01 57.40 21.92 44.86
-716 DMoN–DPR (V) 8.02 57.35 21.97 45.16
-383 DMoN–DPR (V) 12.56 53.64 26.03 49.14
-277 DMoN–DPR (V) 7.78 57.61 21.75 44.89
-274 DMoN–DPR (V) 7.94 57.44 21.59 44.78
-188 DMoN–DPR (V) 7.91 57.49 21.86 45.04
-796 DMoN–DPR (V) 7.92 57.54 21.78 44.83
-Mean 8.66 57.02 22.68 45.18
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (V) 10.94 55.54 22.54 40.25
+550 DMoN–DPR (V) 7.64 58.82 26.91 49.52
+243 DMoN–DPR (V) 7.94 57.61 21.76 44.54
+16 DMoN–DPR (V) 7.96 57.63 21.73 44.37
+716 DMoN–DPR (V) 8.05 57.34 21.70 44.97
+383 DMoN–DPR (V) 9.78 56.13 22.85 37.03
+277 DMoN–DPR (V) 7.80 57.63 21.69 44.89
+274 DMoN–DPR (V) 7.91 57.45 21.60 44.79
+188 DMoN–DPR (V) 10.09 55.60 21.38 37.07
+796 DMoN–DPR (V) 7.98 57.50 21.76 44.86
+Mean $\pm$ Std 8.61 $\pm$ 1.19 57.12 $\pm$ 1.04 22.39 $\pm$ 1.65 43.23 $\pm$ 3.93
 Table 32: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (E) 10.64 55.82 21.91 39.59
-550 DMoN–DPR (E) 7.92 58.29 26.08 48.88
-243 DMoN–DPR (E) 7.88 57.65 21.90 44.67
-16 DMoN–DPR (E) 8.01 57.40 21.90 44.83
-716 DMoN–DPR (E) 8.03 57.34 21.96 45.13
-383 DMoN–DPR (E) 12.55 53.65 26.04 49.15
-277 DMoN–DPR (E) 7.78 57.61 21.75 44.91
-274 DMoN–DPR (E) 7.96 57.42 21.63 44.80
-188 DMoN–DPR (E) 7.93 57.47 21.88 45.06
-796 DMoN–DPR (E) 7.95 57.52 21.79 44.85
-Mean 8.67 57.02 22.68 45.19
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (E) 10.96 55.52 22.51 40.19
+550 DMoN–DPR (E) 7.65 58.81 26.90 49.50
+243 DMoN–DPR (E) 7.94 57.61 21.76 44.55
+16 DMoN–DPR (E) 7.97 57.62 21.74 44.37
+716 DMoN–DPR (E) 8.05 57.33 21.63 44.91
+383 DMoN–DPR (E) 9.82 56.08 22.79 36.95
+277 DMoN–DPR (E) 7.80 57.63 21.69 44.89
+274 DMoN–DPR (E) 7.91 57.45 21.60 44.80
+188 DMoN–DPR (E) 10.08 55.56 21.26 37.16
+796 DMoN–DPR (E) 7.98 57.51 21.75 44.85
+Mean $\pm$ Std 8.62 $\pm$ 1.19 57.11 $\pm$ 1.05 22.36 $\pm$ 1.66 43.22 $\pm$ 3.92
 Table 33: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and variance weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DV) 10.63 55.84 21.89 39.58
-550 DMoN–DPR (DV) 7.93 58.27 26.15 48.92
-243 DMoN–DPR (DV) 7.89 57.64 21.94 44.67
-16 DMoN–DPR (DV) 7.97 57.43 21.83 44.78
-716 DMoN–DPR (DV) 8.06 57.31 22.00 45.11
-383 DMoN–DPR (DV) 12.52 53.69 25.81 49.01
-277 DMoN–DPR (DV) 7.79 57.59 21.72 44.91
-274 DMoN–DPR (DV) 7.93 57.46 21.73 44.92
-188 DMoN–DPR (DV) 7.91 57.48 21.83 45.03
-796 DMoN–DPR (DV) 7.93 57.53 21.86 44.93
-Mean 8.66 57.02 22.68 45.19
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (DV) 10.84 55.65 22.38 40.02
+550 DMoN–DPR (DV) 7.65 58.81 26.87 49.55
+243 DMoN–DPR (DV) 7.95 57.60 21.79 44.56
+16 DMoN–DPR (DV) 7.99 57.59 21.74 44.37
+716 DMoN–DPR (DV) 8.06 57.35 21.72 44.96
+383 DMoN–DPR (DV) 9.76 56.15 22.84 37.08
+277 DMoN–DPR (DV) 7.81 57.62 21.67 44.88
+274 DMoN–DPR (DV) 7.82 57.53 21.67 44.91
+188 DMoN–DPR (DV) 10.12 55.55 21.30 37.10
+796 DMoN–DPR (DV) 7.97 57.50 21.79 44.92
+Mean $\pm$ Std 8.60 $\pm$ 1.17 57.13 $\pm$ 1.03 22.38 $\pm$ 1.64 43.23 $\pm$ 3.95
 Table 34: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, and entropy weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DE) 10.60 55.87 21.90 39.66
-550 DMoN–DPR (DE) 7.94 58.27 26.15 48.91
-243 DMoN–DPR (DE) 7.89 57.64 21.93 44.67
-16 DMoN–DPR (DE) 7.97 57.43 21.83 44.78
-716 DMoN–DPR (DE) 8.08 57.29 22.01 45.14
-383 DMoN–DPR (DE) 12.54 53.68 25.80 49.00
-277 DMoN–DPR (DE) 7.79 57.59 21.71 44.90
-274 DMoN–DPR (DE) 7.97 57.42 21.75 44.94
-188 DMoN–DPR (DE) 7.93 57.47 21.86 45.01
-796 DMoN–DPR (DE) 7.92 57.54 21.78 44.87
-Mean 8.66 57.02 22.67 45.19
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (DE) 10.83 55.65 22.42 40.03
+550 DMoN–DPR (DE) 7.65 58.80 26.87 49.55
+243 DMoN–DPR (DE) 7.95 57.60 21.79 44.57
+16 DMoN–DPR (DE) 8.02 57.57 21.69 44.34
+716 DMoN–DPR (DE) 8.03 57.37 21.77 45.03
+383 DMoN–DPR (DE) 9.81 56.10 22.81 37.07
+277 DMoN–DPR (DE) 7.81 57.62 21.67 44.87
+274 DMoN–DPR (DE) 7.82 57.52 21.68 44.92
+188 DMoN–DPR (DE) 10.10 55.50 21.32 37.24
+796 DMoN–DPR (DE) 7.98 57.50 21.79 44.92
+Mean $\pm$ Std 8.60 $\pm$ 1.17 57.12 $\pm$ 1.04 22.38 $\pm$ 1.63 43.25 $\pm$ 3.93
 Table 35: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (VE) 10.64 55.83 21.89 39.57
-550 DMoN–DPR (VE) 7.92 58.29 26.09 48.88
-243 DMoN–DPR (VE) 7.87 57.65 21.89 44.67
-16 DMoN–DPR (VE) 8.01 57.40 21.91 44.85
-716 DMoN–DPR (VE) 8.03 57.33 21.96 45.13
-383 DMoN–DPR (VE) 12.57 53.63 26.02 49.14
-277 DMoN–DPR (VE) 7.78 57.60 21.74 44.91
-274 DMoN–DPR (VE) 7.96 57.42 21.61 44.81
-188 DMoN–DPR (VE) 7.92 57.47 21.90 45.03
-796 DMoN–DPR (VE) 7.94 57.52 21.77 44.84
-Mean 8.66 57.01 22.68 45.18
-Variance 0.03 0.02 0.03 0.07
+993 DMoN–DPR (VE) 10.85 55.64 22.48 40.10
+550 DMoN–DPR (VE) 7.65 58.82 26.90 49.50
+243 DMoN–DPR (VE) 7.94 57.61 21.76 44.54
+16 DMoN–DPR (VE) 7.97 57.62 21.74 44.37
+716 DMoN–DPR (VE) 8.06 57.33 21.63 44.91
+383 DMoN–DPR (VE) 9.80 56.09 22.79 36.98
+277 DMoN–DPR (VE) 7.80 57.63 21.68 44.88
+274 DMoN–DPR (VE) 7.92 57.44 21.60 44.79
+188 DMoN–DPR (VE) 10.01 55.65 21.43 37.15
+796 DMoN–DPR (VE) 7.97 57.50 21.73 44.84
+Mean $\pm$ Std 8.60 $\pm$ 1.16 57.13 $\pm$ 1.02 22.37 $\pm$ 1.65 43.21 $\pm$ 3.93
 Table 36: Results obtained by using DMoN–DPR pooling using the best epsilon value of 0.001, distance weight of 1, variance weight of 0.001, and entropy weight of 0.001 on the PubMed dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DVE) 10.59 55.87 21.90 39.65
-550 DMoN–DPR (DVE) 7.93 58.27 26.14 48.91
-243 DMoN–DPR (DVE) 7.89 57.64 21.93 44.67
-16 DMoN–DPR (DVE) 7.97 57.42 21.84 44.78
-716 DMoN–DPR (DVE) 8.08 57.29 21.91 45.06
-383 DMoN–DPR (DVE) 12.51 53.69 25.79 48.98
-277 DMoN–DPR (DVE) 7.79 57.58 21.70 44.88
-274 DMoN–DPR (DVE) 7.94 57.45 21.74 44.94
-188 DMoN–DPR (DVE) 7.94 57.45 21.89 45.04
-796 DMoN–DPR (DVE) 7.92 57.54 21.81 44.89
-Mean 8.66 57.02 22.67 45.18
-Variance 0.03 0.02 0.03 0.07
-B.4 Coauthor CS
-The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
+993 DMoN–DPR (DVE) 10.85 55.63 22.40 40.04
+550 DMoN–DPR (DVE) 7.66 58.80 26.86 49.54
+243 DMoN–DPR (DVE) 7.94 57.61 21.79 44.57
+16 DMoN–DPR (DVE) 8.02 57.57 21.69 44.34
+716 DMoN–DPR (DVE) 8.07 57.34 21.68 44.93
+383 DMoN–DPR (DVE) 9.81 56.09 22.81 37.08
+277 DMoN–DPR (DVE) 7.81 57.62 21.67 44.87
+274 DMoN–DPR (DVE) 7.83 57.51 21.69 44.93
+188 DMoN–DPR (DVE) 10.15 55.47 21.31 37.20
+796 DMoN–DPR (DVE) 7.97 57.50 21.75 44.87
+Mean $\pm$ Std 8.61 $\pm$ 1.18 57.11 $\pm$ 1.05 22.37 $\pm$ 1.64 43.24 $\pm$ 3.92
+E.4 Coauthor CS
+The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the Coauthor CS dataset.
+Table 37: Results obtained by using DiffPool pooling on the Coauthor CS dataset.
+Seed Method Conductance Modularity NMI F1
+993 DiffPool 16.20 61.90 53.06 53.73
+550 DiffPool 19.88 62.52 54.40 46.18
+243 DiffPool 20.41 59.39 49.01 49.43
+16 DiffPool 16.41 62.75 53.63 56.78
+716 DiffPool 16.64 61.46 54.03 57.47
+383 DiffPool 19.21 62.93 54.93 56.09
+277 DiffPool 20.95 63.84 56.25 47.35
+274 DiffPool 16.04 62.92 53.34 52.78
+188 DiffPool 20.58 61.21 56.02 56.85
+796 DiffPool 15.61 63.44 54.53 58.08
+Mean $\pm$ Std 18.19 $\pm$ 2.18 62.24 $\pm$ 1.30 53.92 $\pm$ 2.02 53.47 $\pm$ 4.40
+Table 38: Results obtained by using MinCut pooling on the Coauthor CS dataset.
+Seed Method Conductance Modularity NMI F1
+993 MinCut 22.08 70.93 63.96 46.60
+550 MinCut 21.02 71.89 64.48 47.58
+243 MinCut 21.43 71.52 63.83 49.56
+16 MinCut 21.38 71.57 63.92 49.60
+716 MinCut 21.53 71.33 61.94 47.48
+383 MinCut 21.26 71.68 65.02 50.28
+277 MinCut 21.09 71.80 65.17 49.75
+274 MinCut 21.33 71.64 65.19 49.55
+188 MinCut 21.39 71.49 64.73 50.51
+796 MinCut 21.05 71.92 65.03 49.13
+Mean $\pm$ Std 21.36 $\pm$ 0.31 71.58 $\pm$ 0.29 64.33 $\pm$ 0.99 49.00 $\pm$ 1.31
 Table 39: Results obtained by using DMoN pooling on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN 19.07 72.23 68.48 58.27
-550 DMoN 19.14 71.64 66.78 55.19
-243 DMoN 18.11 72.88 70.78 60.88
-16 DMoN 18.28 73.04 71.13 62.23
-716 DMoN 18.41 72.63 69.20 59.44
-383 DMoN 18.45 72.43 69.96 59.60
-277 DMoN 19.53 72.25 67.51 54.79
-274 DMoN 20.20 71.69 69.38 58.09
-188 DMoN 17.89 73.05 71.52 60.61
-796 DMoN 18.89 72.50 69.77 59.18
-Mean 18.80 72.43 69.45 58.83
-Variance 0.01 0.00 0.02 0.06
+993 DMoN 18.28 72.47 68.53 58.97
+550 DMoN 18.01 72.56 67.90 57.26
+243 DMoN 18.50 72.80 70.05 59.31
+16 DMoN 18.04 72.93 70.70 62.98
+716 DMoN 18.23 72.24 68.54 62.21
+383 DMoN 17.72 73.06 70.11 60.39
+277 DMoN 20.05 72.04 66.73 54.74
+274 DMoN 18.13 73.24 71.11 62.35
+188 DMoN 18.49 73.22 71.11 60.66
+796 DMoN 20.81 71.42 67.82 53.72
+Mean $\pm$ Std 18.63 $\pm$ 0.99 72.60 $\pm$ 0.58 69.26 $\pm$ 1.55 59.26 $\pm$ 3.17
 Table 40: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, and distance weight of 1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR(D) 18.88 72.53 72.11 62.71
-550 DMoN–DPR(D) 20.80 71.19 68.18 57.76
-243 DMoN–DPR(D) 19.36 71.83 71.19 62.87
-16 DMoN–DPR(D) 18.99 72.42 71.12 63.30
-716 DMoN–DPR(D) 17.85 73.08 70.14 56.54
-383 DMoN–DPR(D) 19.24 71.60 69.23 60.49
-277 DMoN–DPR(D) 18.20 73.02 72.35 64.04
-274 DMoN–DPR(D) 18.86 72.11 70.97 61.56
-188 DMoN–DPR(D) 19.85 71.18 68.61 59.59
-796 DMoN–DPR(D) 18.02 72.72 71.70 62.08
-Mean 19.01 72.17 70.56 61.09
-Variance 0.01 0.00 0.02 0.06
+993 DMoN–DPR (D) 19.52 71.62 70.00 60.13
+550 DMoN–DPR (D) 18.63 72.19 70.76 61.55
+243 DMoN–DPR (D) 19.37 71.83 69.94 60.27
+16 DMoN–DPR (D) 19.14 72.53 72.67 64.33
+716 DMoN–DPR (D) 18.83 72.46 70.96 62.26
+383 DMoN–DPR (D) 19.05 72.16 71.26 62.23
+277 DMoN–DPR (D) 18.70 72.93 72.60 63.38
+274 DMoN–DPR (D) 18.46 72.51 71.16 61.65
+188 DMoN–DPR (D) 19.29 71.97 71.58 61.29
+796 DMoN–DPR (D) 17.94 72.77 70.78 61.08
+Mean $\pm$ Std 18.89 $\pm$ 0.48 72.30 $\pm$ 0.42 71.17 $\pm$ 0.93 61.82 $\pm$ 1.30
 Table 41: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (V) 18.87 72.35 68.79 58.77
-550 DMoN–DPR (V) 19.03 71.68 67.07 56.15
-243 DMoN–DPR (V) 18.27 72.74 71.13 61.15
-16 DMoN–DPR (V) 18.28 72.94 71.38 62.48
-716 DMoN–DPR (V) 19.77 71.86 68.11 57.05
-383 DMoN–DPR (V) 18.53 72.22 70.15 60.21
-277 DMoN–DPR (V) 19.78 72.14 67.97 55.30
-274 DMoN–DPR (V) 19.57 72.49 70.45 59.36
-188 DMoN–DPR (V) 18.24 72.89 71.02 59.87
-796 DMoN–DPR (V) 18.52 72.76 71.57 60.67
-Mean 18.89 72.41 69.76 59.10
-Variance 0.00 0.00 0.03 0.05
+993 DMoN–DPR (V) 18.09 72.79 69.54 60.69
+550 DMoN–DPR (V) 19.20 72.05 69.23 60.66
+243 DMoN–DPR (V) 17.60 73.27 71.19 62.29
+16 DMoN–DPR (V) 18.45 72.43 68.76 60.54
+716 DMoN–DPR (V) 18.58 72.35 68.53 59.02
+383 DMoN–DPR (V) 19.05 72.25 69.61 59.12
+277 DMoN–DPR (V) 19.33 72.27 69.12 57.74
+274 DMoN–DPR (V) 18.62 72.97 70.38 62.66
+188 DMoN–DPR (V) 18.55 73.04 71.22 60.91
+796 DMoN–DPR (V) 20.01 71.56 67.98 55.77
+Mean $\pm$ Std 18.75 $\pm$ 0.68 72.50 $\pm$ 0.52 69.56 $\pm$ 1.08 59.94 $\pm$ 2.08
 Table 42: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (E) 18.95 71.61 72.94 59.97
-550 DMoN–DPR (E) 19.87 71.23 70.05 56.24
-243 DMoN–DPR (E) 20.13 70.45 71.39 60.47
-16 DMoN–DPR (E) 19.44 71.06 72.47 64.02
-716 DMoN–DPR (E) 19.11 71.28 73.55 64.42
-383 DMoN–DPR (E) 20.09 70.20 67.61 54.84
-277 DMoN–DPR (E) 21.27 69.71 68.97 56.84
-274 DMoN–DPR (E) 20.01 71.45 73.13 62.46
-188 DMoN–DPR (E) 18.70 71.68 72.18 63.95
-796 DMoN–DPR (E) 19.74 71.04 71.06 60.77
-Mean 19.73 70.97 71.34 60.40
-Variance 0.01 0.00 0.04 0.12
+993 DMoN–DPR (E) 20.73 69.18 70.29 57.08
+550 DMoN–DPR (E) 19.82 71.30 70.48 57.63
+243 DMoN–DPR (E) 19.35 71.39 72.46 64.00
+16 DMoN–DPR (E) 18.71 71.88 73.81 66.15
+716 DMoN–DPR (E) 18.77 72.20 74.36 66.49
+383 DMoN–DPR (E) 19.01 71.17 70.30 60.87
+277 DMoN–DPR (E) 21.91 69.36 70.12 59.57
+274 DMoN–DPR (E) 19.90 71.03 72.03 60.58
+188 DMoN–DPR (E) 19.49 71.84 73.02 62.90
+796 DMoN–DPR (E) 20.85 69.82 68.94 58.00
+Mean $\pm$ Std 19.85 $\pm$ 1.03 70.92 $\pm$ 1.08 71.58 $\pm$ 1.81 61.33 $\pm$ 3.43
 Table 43: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and variance weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DV) 18.76 72.48 72.77 63.41
-550 DMoN–DPR (DV) 19.71 71.95 70.36 61.52
-243 DMoN–DPR (DV) 19.10 72.09 71.75 63.65
-16 DMoN–DPR (DV) 18.57 72.77 73.03 63.98
-716 DMoN–DPR (DV) 19.02 72.54 70.44 58.08
-383 DMoN–DPR (DV) 18.57 72.33 72.29 63.93
-277 DMoN–DPR (DV) 17.94 73.14 72.46 63.13
-274 DMoN–DPR (DV) 19.41 71.73 70.41 60.45
-188 DMoN–DPR (DV) 20.33 71.18 69.03 58.99
-796 DMoN–DPR (DV) 18.06 72.38 70.79 60.99
-Mean 18.95 72.26 71.33 61.81
-Variance 0.01 0.00 0.02 0.05
+993 DMoN–DPR (DV) 19.55 71.66 69.75 60.56
+550 DMoN–DPR (DV) 18.24 72.37 71.14 61.70
+243 DMoN–DPR (DV) 18.94 71.86 70.04 60.62
+16 DMoN–DPR (DV) 19.75 71.81 72.52 63.87
+716 DMoN–DPR (DV) 20.11 71.14 68.74 58.85
+383 DMoN–DPR (DV) 19.52 71.72 69.75 60.47
+277 DMoN–DPR (DV) 18.63 72.57 72.00 62.98
+274 DMoN–DPR (DV) 18.68 72.21 71.43 62.77
+188 DMoN–DPR (DV) 19.89 71.46 69.60 58.73
+796 DMoN–DPR (DV) 18.13 72.81 72.20 62.90
+Mean $\pm$ Std 19.14 $\pm$ 0.71 71.96 $\pm$ 0.52 70.72 $\pm$ 1.30 61.35 $\pm$ 1.78
 Table 44: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, and entropy weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DE) 19.82 70.82 71.87 64.34
-550 DMoN–DPR (DE) 19.72 70.51 70.06 61.80
-243 DMoN–DPR (DE) 19.74 70.31 70.33 61.84
-16 DMoN–DPR (DE) 19.83 71.73 74.31 63.08
-716 DMoN–DPR (DE) 18.97 71.30 71.50 62.63
-383 DMoN–DPR (DE) 19.97 70.74 72.47 65.26
-277 DMoN–DPR (DE) 19.50 71.15 72.02 60.54
-274 DMoN–DPR (DE) 19.42 69.84 68.51 59.93
-188 DMoN–DPR (DE) 21.62 69.25 67.61 56.77
-796 DMoN–DPR (DE) 20.19 69.93 69.69 61.41
-Mean 19.88 70.56 70.84 61.76
-Variance 0.00 0.01 0.04 0.06
+993 DMoN–DPR (DE) 19.75 70.62 71.29 63.49
+550 DMoN–DPR (DE) 20.49 70.46 71.58 63.84
+243 DMoN–DPR (DE) 20.83 69.25 69.19 60.23
+16 DMoN–DPR (DE) 19.57 70.68 70.86 62.53
+716 DMoN–DPR (DE) 19.12 71.23 71.83 62.71
+383 DMoN–DPR (DE) 19.22 71.16 72.81 64.86
+277 DMoN–DPR (DE) 19.78 71.78 73.49 65.51
+274 DMoN–DPR (DE) 19.52 70.15 70.28 61.17
+188 DMoN–DPR (DE) 20.82 69.37 67.33 57.40
+796 DMoN–DPR (DE) 19.00 70.96 70.94 61.83
+Mean $\pm$ Std 19.81 $\pm$ 0.68 70.57 $\pm$ 0.80 70.96 $\pm$ 1.76 62.36 $\pm$ 2.37
 Table 45: Results obtained by using DMoN–DPR pooling using the best variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (VE) 19.08 71.49 73.11 60.17
-550 DMoN–DPR (VE) 19.95 71.16 70.08 56.20
-243 DMoN–DPR (VE) 20.21 70.35 71.26 60.33
-16 DMoN–DPR (VE) 19.62 70.89 72.13 63.50
-716 DMoN–DPR (VE) 19.16 71.25 73.42 64.30
-383 DMoN–DPR (VE) 20.23 70.06 67.62 54.87
-277 DMoN–DPR (VE) 20.14 70.81 70.43 59.06
-274 DMoN–DPR (VE) 20.04 71.40 73.23 62.31
-188 DMoN–DPR (VE) 18.72 71.67 72.17 63.93
-796 DMoN–DPR (VE) 19.80 70.97 71.09 60.92
-Mean 19.70 71.01 71.45 60.56
-Variance 0.00 0.00 0.03 0.10
+993 DMoN–DPR (VE) 20.74 69.08 70.35 57.99
+550 DMoN–DPR (VE) 19.85 71.25 70.50 57.59
+243 DMoN–DPR (VE) 19.43 71.30 72.55 64.08
+16 DMoN–DPR (VE) 18.74 71.85 73.81 66.17
+716 DMoN–DPR (VE) 19.22 71.37 73.19 65.61
+383 DMoN–DPR (VE) 18.65 71.43 70.49 60.75
+277 DMoN–DPR (VE) 21.92 69.32 70.10 59.50
+274 DMoN–DPR (VE) 19.93 70.96 71.96 60.58
+188 DMoN–DPR (VE) 19.44 71.87 73.01 63.01
+796 DMoN–DPR (VE) 20.99 69.70 68.72 57.97
+Mean $\pm$ Std 19.89 $\pm$ 1.04 70.81 $\pm$ 1.04 71.47 $\pm$ 1.66 61.33 $\pm$ 3.21
 Table 46: Results obtained by using DMoN–DPR pooling using the best epsilon value of 1, distance weight of 1, variance weight of 0.1, and entropy weight of 0.1 on the Coauthor CS dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DVE) 19.32 71.13 72.76 65.12
-550 DMoN–DPR (DVE) 19.75 70.50 70.00 61.78
-243 DMoN–DPR (DVE) 19.97 70.19 70.40 61.86
-16 DMoN–DPR (DVE) 18.87 72.23 74.23 65.17
-716 DMoN–DPR (DVE) 19.03 71.26 71.56 62.76
-383 DMoN–DPR (DVE) 20.06 70.65 72.47 65.18
-277 DMoN–DPR (DVE) 19.06 72.07 73.97 64.36
-274 DMoN–DPR (DVE) 19.41 69.81 68.57 60.03
-188 DMoN–DPR (DVE) 21.66 69.20 67.65 56.86
-796 DMoN–DPR (DVE) 20.30 69.78 69.21 60.98
-Mean 19.74 70.68 71.08 62.41
-Variance 0.01 0.01 0.05 0.07
-B.5 Coauthor Physics
-The following tables list the performance of DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.
+993 DMoN–DPR (DVE) 20.34 70.31 70.86 63.39
+550 DMoN–DPR (DVE) 20.09 70.77 71.87 64.16
+243 DMoN–DPR (DVE) 19.56 70.27 70.03 60.73
+16 DMoN–DPR (DVE) 20.21 70.65 72.33 63.54
+716 DMoN–DPR (DVE) 19.26 71.08 71.81 62.97
+383 DMoN–DPR (DVE) 19.17 71.20 72.88 64.81
+277 DMoN–DPR (DVE) 18.96 72.10 73.54 65.43
+274 DMoN–DPR (DVE) 19.65 70.36 70.79 61.93
+188 DMoN–DPR (DVE) 20.98 69.20 67.10 57.05
+796 DMoN–DPR (DVE) 18.70 71.14 71.62 62.70
+Mean $\pm$ Std 19.69 $\pm$ 0.71 70.71 $\pm$ 0.76 71.28 $\pm$ 1.79 62.67 $\pm$ 2.40
+E.5 Coauthor Physics
+The following tables list the performance of DiffPool, MinCut pooling, DMoN and different DMoN–DPR configurations across different seeds on the Coauthor Physics dataset.
+Table 47: Results obtained by using DiffPool pooling on the Coauthor Physics dataset.
+Seed Method Conductance Modularity NMI F1
+993 DiffPool 13.06 56.73 57.94 53.93
+550 DiffPool 15.25 59.42 59.33 53.71
+243 DiffPool 13.06 57.18 56.61 52.20
+16 DiffPool 13.26 56.55 57.69 53.75
+716 DiffPool 14.42 58.97 61.09 53.65
+383 DiffPool 14.03 57.15 58.03 54.06
+277 DiffPool 13.37 56.48 47.24 46.20
+274 DiffPool 13.26 55.70 58.06 56.79
+188 DiffPool 12.23 57.04 58.30 54.31
+796 DiffPool 17.18 59.19 47.89 38.47
+Mean $\pm$ Std 13.91 $\pm$ 1.42 57.44 $\pm$ 1.29 56.22 $\pm$ 4.71 51.71 $\pm$ 5.38
+Table 48: Results obtained by using MinCut pooling on the Coauthor Physics dataset.
+Seed Method Conductance Modularity NMI F1
+993 MinCut 14.08 61.78 48.91 45.03
+550 MinCut 14.35 61.54 56.89 51.00
+243 MinCut 13.74 61.79 51.41 47.76
+16 MinCut 13.38 62.16 57.19 51.39
+716 MinCut 13.18 62.55 57.10 51.27
+383 MinCut 13.95 62.12 48.62 42.59
+277 MinCut 14.17 61.39 48.28 46.92
+274 MinCut 14.20 61.31 45.80 46.44
+188 MinCut 13.33 62.51 52.14 47.82
+796 MinCut 14.95 60.38 47.53 47.70
+Mean $\pm$ Std 13.93 $\pm$ 0.54 61.75 $\pm$ 0.65 51.39 $\pm$ 4.30 47.79 $\pm$ 2.84
 Table 49: Results obtained by using DMoN pooling on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN 12.61 60.14 44.16 46.45
-550 DMoN 14.36 63.58 51.07 45.68
-243 DMoN 13.84 63.82 56.68 49.40
-16 DMoN 13.81 63.78 57.00 49.64
-716 DMoN 13.91 63.70 55.59 49.16
-383 DMoN 13.84 63.08 48.79 45.27
-277 DMoN 14.06 63.70 46.72 39.63
-274 DMoN 13.84 63.93 53.70 46.41
-188 DMoN 13.94 63.70 55.66 49.31
-796 DMoN 13.89 64.04 54.63 47.03
-Mean 13.81 63.35 52.40 46.80
-Variance 0.00 0.01 0.20 0.09
+993 DMoN 12.64 59.92 44.49 46.04
+550 DMoN 13.87 64.01 54.36 47.54
+243 DMoN 13.89 63.77 56.67 49.51
+16 DMoN 13.82 63.77 56.89 49.55
+716 DMoN 13.93 63.72 55.42 49.02
+383 DMoN 13.65 63.63 52.66 46.65
+277 DMoN 13.48 63.85 50.84 44.74
+274 DMoN 13.84 64.02 53.53 46.36
+188 DMoN 13.89 63.73 55.87 49.33
+796 DMoN 14.02 64.12 54.29 46.36
+Mean $\pm$ Std 13.70 $\pm$ 0.40 63.45 $\pm$ 1.25 53.50 $\pm$ 3.67 47.51 $\pm$ 1.73
 Table 50: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, and distance weight of 1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (D) 16.08 60.64 49.75 44.05
-550 DMoN–DPR (D) 15.42 56.02 50.59 52.21
-243 DMoN–DPR (D) 15.78 58.21 58.58 53.21
-16 DMoN–DPR (D) 15.91 54.69 53.21 54.59
-716 DMoN–DPR (D) 15.85 57.35 50.14 49.28
-383 DMoN–DPR (D) 18.82 55.05 51.96 48.46
-277 DMoN–DPR (D) 19.89 54.78 50.10 47.08
-274 DMoN–DPR (D) 15.03 56.06 55.73 54.53
-188 DMoN–DPR (D) 13.71 59.75 61.71 54.99
-796 DMoN–DPR (D) 18.02 56.74 58.17 51.22
-Mean 16.45 56.93 53.99 50.96
-Variance 0.04 0.04 0.18 0.13
+993 DMoN–DPR (D) 14.00 62.17 51.59 47.84
+550 DMoN–DPR (D) 19.30 56.61 50.82 46.32
+243 DMoN–DPR (D) 18.21 57.77 58.61 50.55
+16 DMoN–DPR (D) 16.19 54.39 53.19 55.67
+716 DMoN–DPR (D) 16.93 55.94 48.09 49.54
+383 DMoN–DPR (D) 19.58 55.58 52.71 46.44
+277 DMoN–DPR (D) 14.97 55.73 51.49 53.53
+274 DMoN–DPR (D) 15.55 56.74 59.80 55.95
+188 DMoN–DPR (D) 14.85 55.96 55.95 54.65
+796 DMoN–DPR (D) 14.78 58.04 52.68 49.37
+Mean $\pm$ Std 16.44 $\pm$ 1.99 56.89 $\pm$ 2.14 53.49 $\pm$ 3.61 50.99 $\pm$ 3.71
 Table 51: Results obtained by using DMoN–DPR pooling using the best variance weight of 1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (V) 12.00 58.02 45.50 45.54
-550 DMoN–DPR (V) 12.44 57.44 53.99 54.11
-243 DMoN–DPR (V) 13.82 61.60 59.00 51.98
-16 DMoN–DPR (V) 13.64 58.69 58.90 56.75
-716 DMoN–DPR (V) 14.06 61.38 56.03 51.89
-383 DMoN–DPR (V) 14.42 63.20 52.72 45.15
-277 DMoN–DPR (V) 12.96 57.38 52.20 53.30
-274 DMoN–DPR (V) 12.88 57.23 54.09 56.20
-188 DMoN–DPR (V) 13.95 57.14 52.27 54.06
-796 DMoN–DPR (V) 14.29 61.71 55.55 49.94
-Mean 13.45 59.38 54.03 51.89
-Variance 0.01 0.05 0.15 0.16
+993 DMoN–DPR (V) 12.28 58.05 45.73 45.50
+550 DMoN–DPR (V) 11.94 57.38 54.27 54.78
+243 DMoN–DPR (V) 13.49 60.85 59.11 53.04
+16 DMoN–DPR (V) 13.55 58.65 58.50 56.62
+716 DMoN–DPR (V) 11.83 58.84 56.84 53.74
+383 DMoN–DPR (V) 14.44 57.84 54.21 55.65
+277 DMoN–DPR (V) 12.03 58.71 57.09 53.73
+274 DMoN–DPR (V) 14.15 57.02 53.90 55.44
+188 DMoN–DPR (V) 13.11 57.40 54.62 53.97
+796 DMoN–DPR (V) 15.02 61.14 45.01 45.34
+Mean $\pm$ Std 13.18 $\pm$ 1.14 58.59 $\pm$ 1.41 53.93 $\pm$ 4.87 52.78 $\pm$ 4.02
 Table 52: Results obtained by using DMoN–DPR pooling using the best entropy weight of 0.1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (E) 12.88 58.41 44.16 45.67
-550 DMoN–DPR (E) 12.07 57.62 54.04 53.83
-243 DMoN–DPR (E) 13.79 62.02 58.66 51.38
-16 DMoN–DPR (E) 13.77 60.22 58.66 53.25
-716 DMoN–DPR (E) 14.41 61.97 56.14 50.58
-383 DMoN–DPR (E) 13.88 58.47 53.74 53.85
-277 DMoN–DPR (E) 12.85 57.66 52.26 52.98
-274 DMoN–DPR (E) 15.10 59.77 51.39 45.72
-188 DMoN–DPR (E) 14.12 62.20 56.74 50.19
-796 DMoN–DPR (E) 15.15 62.10 45.87 43.09
-Mean 13.80 60.04 53.17 50.05
-Variance 0.01 0.04 0.25 0.15
+993 DMoN–DPR (E) 13.59 59.64 48.13 45.60
+550 DMoN–DPR (E) 11.94 57.60 54.02 54.19
+243 DMoN–DPR (E) 13.72 61.90 58.72 51.62
+16 DMoN–DPR (E) 13.53 58.94 57.15 55.16
+716 DMoN–DPR (E) 14.17 57.80 47.71 51.12
+383 DMoN–DPR (E) 14.24 58.41 53.68 54.15
+277 DMoN–DPR (E) 11.57 58.08 53.16 52.07
+274 DMoN–DPR (E) 13.26 57.99 53.77 51.77
+188 DMoN–DPR (E) 14.08 60.75 55.54 50.03
+796 DMoN–DPR (E) 14.81 62.19 46.39 45.17
+Mean $\pm$ Std 13.49 $\pm$ 1.02 59.33 $\pm$ 1.72 52.83 $\pm$ 4.13 51.09 $\pm$ 3.39
 Table 53: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and variance weight 1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DV) 14.08 60.60 47.53 46.00
-550 DMoN–DPR (DV) 14.07 55.94 54.18 57.79
-243 DMoN–DPR (DV) 13.03 58.15 60.48 59.90
-16 DMoN–DPR (DV) 12.03 55.51 54.29 57.77
-716 DMoN–DPR (DV) 12.54 56.50 52.40 57.28
-383 DMoN–DPR (DV) 13.82 56.10 52.59 57.34
-277 DMoN–DPR (DV) 13.96 56.84 53.28 57.82
-274 DMoN–DPR (DV) 12.40 56.41 52.47 57.72
-188 DMoN–DPR (DV) 14.36 56.80 60.08 57.68
-796 DMoN–DPR (DV) 14.34 55.61 53.64 54.26
-Mean 13.46 56.85 54.09 56.36
-Variance 0.01 0.02 0.14 0.15
+993 DMoN–DPR (DV) 15.37 55.95 48.78 53.49
+550 DMoN–DPR (DV) 14.84 55.32 52.13 57.57
+243 DMoN–DPR (DV) 13.82 56.19 56.71 57.69
+16 DMoN–DPR (DV) 11.67 56.14 54.87 57.53
+716 DMoN–DPR (DV) 10.97 57.41 62.02 61.81
+383 DMoN–DPR (DV) 13.45 56.41 54.45 57.87
+277 DMoN–DPR (DV) 14.25 56.26 52.86 57.93
+274 DMoN–DPR (DV) 11.79 56.58 57.37 59.70
+188 DMoN–DPR (DV) 14.17 56.86 60.24 58.29
+796 DMoN–DPR (DV) 13.98 57.70 58.95 58.04
+Mean $\pm$ Std 13.43 $\pm$ 1.46 56.48 $\pm$ 0.70 55.84 $\pm$ 4.02 57.99 $\pm$ 2.06
 Table 54: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DE) 15.49 55.62 48.15 53.66
-550 DMoN–DPR (DE) 15.31 55.21 52.10 56.57
-243 DMoN–DPR (DE) 13.74 56.71 59.84 58.07
-16 DMoN–DPR (DE) 12.99 55.20 53.27 56.69
-716 DMoN–DPR (DE) 14.00 56.38 59.34 58.29
-383 DMoN–DPR (DE) 13.94 56.83 55.09 56.84
-277 DMoN–DPR (DE) 14.97 54.26 54.14 60.85
-274 DMoN–DPR (DE) 12.57 55.85 57.05 59.56
-188 DMoN–DPR (DE) 12.99 56.54 58.60 58.86
-796 DMoN–DPR (DE) 16.16 58.41 49.49 49.99
+993 DMoN–DPR (DE) 13.94 61.00 47.60 46.37
+550 DMoN–DPR (DE) 14.11 56.74 54.67 55.98
+243 DMoN–DPR (DE) 13.89 57.00 60.05 58.60
+16 DMoN–DPR (DE) 13.90 54.44 53.68 57.90
+716 DMoN–DPR (DE) 11.35 57.47 63.18 61.39
+383 DMoN–DPR (DE) 15.41 54.93 51.58 53.92
+277 DMoN–DPR (DE) 15.84 54.43 51.48 57.57
+274 DMoN–DPR (DE) 14.14 57.42 53.66 56.28
+188 DMoN–DPR (DE) 15.33 56.20 56.80 55.88
+796 DMoN–DPR (DE) 13.83 59.24 47.47 46.26
+Mean $\pm$ Std 14.17 $\pm$ 1.24 56.89 $\pm$ 2.09 54.02 $\pm$ 5.00 55.02 $\pm$ 4.99
 Mean 14.22 56.10 54.71 56.94
@@
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (VE) 8.50 36.32 14.32 44.25
-550 DMoN–DPR (VE) 13.12 57.28 52.98 56.87
-243 DMoN–DPR (VE) 14.64 59.71 55.12 51.50
-16 DMoN–DPR (VE) 13.92 58.45 62.37 58.80
-716 DMoN–DPR (VE) 13.75 57.27 52.06 51.47
-383 DMoN–DPR (VE) 14.04 57.77 56.72 58.23
-277 DMoN–DPR (VE) 13.04 57.90 54.22 57.55
-274 DMoN–DPR (VE) 12.10 56.56 51.63 56.60
-188 DMoN–DPR (VE) 10.40 55.47 55.62 62.42
-796 DMoN–DPR (VE) 15.01 60.44 46.37 47.48
-Mean 12.85 55.72 50.14 54.52
-Variance 0.04 0.49 1.75 0.32
+993 DMoN–DPR (VE) 8.51 36.26 14.21 44.28
+550 DMoN–DPR (VE) 13.11 57.01 52.67 57.06
+243 DMoN–DPR (VE) 14.40 59.91 56.09 52.22
+16 DMoN–DPR (VE) 13.97 58.17 62.31 59.21
+716 DMoN–DPR (VE) 13.66 58.24 53.74 53.47
+383 DMoN–DPR (VE) 14.12 57.60 56.52 57.86
+277 DMoN–DPR (VE) 16.66 60.87 48.23 44.09
+274 DMoN–DPR (VE) 12.18 56.56 51.47 56.39
+188 DMoN–DPR (VE) 11.16 56.05 56.50 60.71
+796 DMoN–DPR (VE) 13.32 58.12 47.75 50.50
+Mean $\pm$ Std 13.11 $\pm$ 2.17 55.88 $\pm$ 7.04 49.95 $\pm$ 13.27 53.58 $\pm$ 5.86
 Table 56: Results obtained by using DMoN–DPR pooling using the best epsilon value of 10, distance weight of 1, variance weight of 1, and entropy weight of 0.1 on the Coauthor Physics dataset.
 Seed Method Conductance Modularity NMI F1
-993 DMoN–DPR (DVE) 15.43 55.58 47.68 53.96
-550 DMoN–DPR (DVE) 14.27 55.01 53.99 58.87
-243 DMoN–DPR (DVE) 12.76 56.10 53.71 58.34
-16 DMoN–DPR (DVE) 14.62 53.36 49.11 57.45
-716 DMoN–DPR (DVE) 13.06 54.46 57.73 62.18
-383 DMoN–DPR (DVE) 12.92 55.12 58.39 60.97
-277 DMoN–DPR (DVE) 14.04 55.08 51.68 58.26
-274 DMoN–DPR (DVE) 12.59 55.68 58.34 60.52
-188 DMoN–DPR (DVE) 11.41 57.40 63.84 62.83
-796 DMoN–DPR (DVE) 13.51 54.94 48.67 55.26
-Mean 13.46 55.27 54.31 58.86
-Variance 0.01 0.01 0.27 0.08
-Appendix C Code Implementation
+993 DMoN–DPR (DVE) 14.25 57.43 42.12 46.26
+550 DMoN–DPR (DVE) 14.53 55.66 53.54 57.87
+243 DMoN–DPR (DVE) 12.50 55.96 54.67 59.77
+16 DMoN–DPR (DVE) 13.27 54.52 51.00 57.72
+716 DMoN–DPR (DVE) 12.95 54.76 54.90 58.06
+383 DMoN–DPR (DVE) 12.76 55.00 56.93 60.32
+277 DMoN–DPR (DVE) 14.47 54.87 52.66 59.00
+274 DMoN–DPR (DVE) 12.47 55.39 58.52 61.38
+188 DMoN–DPR (DVE) 11.29 57.23 62.59 62.06
+796 DMoN–DPR (DVE) 9.90 53.89 48.04 57.18
+Mean $\pm$ Std 12.84 $\pm$ 1.45 55.47 $\pm$ 1.14 53.50 $\pm$ 5.67 57.96 $\pm$ 4.42
+Appendix F Implementation Details
 The code was implemented by extending the DMoN implementation in PyTorch Geometric (Fey and Lenssen, 2019), and was trained and evaluated using the evaluation protocols found in the official DMoN repository (Tsitsulin et al., 2023).
+The experiments were run on an A100 GPU with 40GB of memory offered by Google Colab Pro.
+Regarding the runtime analysis results and clustering visualization, the code was run on an Apple M2 Max CPU.